### v1

# Understanding Biases in ChatGPT-based Recommender Systems: Provider Fairness, Temporal Stability, and Recency

Abstract.
This study explores the nuanced capabilities and inherent biases of Recommender Systems using Large Language Models (RecLLMs), with a focus on ChatGPT-based systems. It studies into the contrasting behaviors of generative models and traditional collaborative filtering models in movie recommendations. The research primarily investigates prompt design strategies and their impact on various aspects of recommendation quality, including accuracy, provider fairness, diversity, stability, genre dominance, and temporal freshness (recency).
Our experimental analysis reveals that the introduction of specific system roles and prompt strategies in RecLLMs significantly influences their performance. For instance, role-based prompts enhance fairness and diversity in recommendations, mitigating popularity bias. We find that while GPT-based models do not always match the performance of CF baselines, they exhibit a unique tendency to recommend newer and more diverse movie genres. Notably, GPT-based models tend to recommend more recent films, particularly those released post-2000, and show a preference for genres like ‘Drama’ and ‘Comedy,’ and ‘Romance’ (compared to CF ‘Action’, ‘Adventure’) presumably due to the RecLLMs’ training on varied data sets, which allows them to capture recent trends and discussions more effectively than CF models.
Interestingly, our results demonstrate that the Simple and Chain of Thought (COT) paradigms yield the highest accuracy. These findings imply the potential of combining these strategies with scenarios that favor more recent content, thereby offering a more balanced and up-to-date recommendation experience. This study contributes significantly to the understanding of emerging RecLLMs, particularly in the context of harms and biases within these systems. It highlights the complexities and potential of prompt design in impacting recommendation outcomes and paves the way for future research in optimizing accuracy, fairness, and user experience in such systems.
Recommender Systems, Large Language Models, Bias and Fairness in RS, Movie Recommendation Analysis, Prompt Design Strategies, ChatGPT, Stability and Diversity in Recommendations
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXX
1. Introduction
Recommender systems have increasingly become integral in various domains, offering substantial benefits for stakeholders in two-sided, or broadly, multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019). These systems alleviate user information overload and enhance producer sales and revenue. From a learning viewpoint, recommendation models can generally be classified into two categories: ‘discriminative’ and ‘generative’ models (Deldjoo et al., 2022).
Discriminative models, often considered the traditional approach in recommendation systems, function by categorizing and differentiating to inform decisions about user preferences. For example, in the movie domain, these models can be compared to expert critics that are adept at discerning the distinctive characteristics that separate one type of item (movie) from another, given the audience. In other words, this differentiation is informed by factors such as user historical ratings, genre preferences, and other relevant attribute of items. Mainstream collaborative filtering (CF) recommendation models, such as matrix factorization (MF)-based models and more complex neural variations such as NeuMF (He et al., 2017), operate on this principle. Similarly, graph-based models, such as NGCF (Wang et al., 2019) and LightGCN (He et al., 2020), leverage the relational structure of the data to improve the quality of recommendations. Regardless of the methodological approach, traditional discriminative models are heavily based on knowledge derived from limited datasets and specific domains to calculate recommendations and inform decisions.
Generative models represent a transformative approach, focusing on understanding the underlying structure and distribution of data to generate authentic-looking content. Their application in recommender systems (RS) exemplifies a shift from simple preference matching to a more nuanced understanding of user behavior and interests. In the context of movie recommender systems, these models can be likened to hypothetical curators who not only distinguish between various items but also comprehend the overarching patterns and tastes of the audience. Thus, these models go beyond mere differentiation; they synthesize and anticipate user preferences, providing tailored recommendations.
Large Language Models (LLMs), exemplified by the Generative Pretrained Transformers (GPT), recently heralded a new era in Natural Language Processing (NLP). These Transformer-based models, such as BERT, GPT series, PaLM series, etc., are characterized by their vast number of parameters and are trained on extensive datasets using self/semi-supervised learning techniques. Their training on a wide range of internet text data endows LLMs with the exceptional ability to generate contextually relevant and coherent text in response to input prompts. Different from traditional recommendation systems, LLM-based models excel in capturing contextual information, comprehending user queries, item descriptions, and other textual data more effectively (Geng et al., 2022). Consequently, specifically-designed LLMs for recommender systems (RecLLMs) can harness LLMs to extract high-quality representations of textual features and leverage extensive external knowledge encoded within LLMs (Liu et al., 2023) to offer more nuanced insights into understanding user preference and profile. This is a significant leap forward compared to traditional RS. This aligns well with the overarching goal of delivering personalized, context-aware recommendations that reflect individual tastes and behaviors.
In general, LLMs can contribute to recommender systems in three principal ways (Wu et al., 2023; Li et al., 2023b; Xu et al., 2024). First, LLMs serve as direct recommenders by generating personalized content and suggestions, stemming from their advanced understanding of context and user queries (Geng et al., 2022; Wu et al., 2023). Second, they enhance traditional models by extracting rich semantic representations from textual data. This aspect is particularly important in addressing the data sparsity issue in historical interactions (Da’u and Salim, 2020). Third, LLMs act as complex simulators in the recommendation process, effectively testing and refining the system’s predictions (Liu et al., 2023). The current work in hand primarily concentrates on the first application, that is using LLMs as recommender, with ChatGPT serving as the case study.
The primary aim of this research is to enhance our understanding of “Provider Fairness within ChatGPT-based RecLLMs”. We focus on uncovering and understanding a range of biases – mostly underexplored in existing literature –, and examining how these biases manifest in comparison to traditional Collaborative Filtering (CF) models. This study looks into the dynamics of these biases, providing a detailed contrast between the two types of systems, and thereby contributing to a more nuanced understanding of the complexities inherent in modern RS.
In particular, this research diligently identifies and analyzes inherent biases inherent in these two types of systems, such as (i) the preference for popular movies, (ii) the balance between old and recent movie recommendations, (iii) the tendency to recommend certain movie genres, (iv) the consistency/stability of recommendations over time. For instance, GPT-based models were found to have a penchant for suggesting recent films like those released post-2000, exemplified by their frequent recommendations of newer titles such as ”Inception” (2010) and ”The Dark Knight” (2008), compared to CF models that often favored classics like ”The Godfather” (1972) and ”Casablanca” (1942). Furthermore, genre dominance was evident, with GPT models showing a distinct inclination towards ‘Drama’ and ‘Comedy’, whereas CF models predominantly recommended ‘Action’ and ‘Sci-Fi’ films. Another crucial aspect of this research is exploring the impact of prompt engineering in a zero-shot learning setting on the performance of RecLLMs in terms of personalization, diversity, recency, and fairness. Zero-shot learning refers to the model’s capacity to handle tasks for which it has not been explicitly trained, relying instead on its general language understanding and adaptability.
Here our study seeks to provide an answer to the following question: Can prompts serve as a controllability tool to steer the recommendation process in a specific direction (increase accuracy, reduce provider unfairness) based on the interests of providers or other stakeholders, and what potential harms are associated with this approach?
Figure 1. Conceptual idea behind our work
Example. We illustrate the impact of different “prompt design” strategies on the output of ChatGPT-based recommendation systems through a clear example, as shown in Figure 1. We use a randomly chosen user profile from the MovieLens dataset to demonstrate this effect. The user’s historical interactions with the system include a diverse range of movies such as “Toy Story” (categorized with genres Adventure—Animation—Children—Comedy—Fantasy), “If Lucy Fell” (Comedy—Romance), and “Hard Target” (Action—Adventure—Crime—Thriller).
To highlight the differences in recommendations based on the design of the system, we designed a total of seven prompts categorized into three distinct classes. These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the system’s outputs can vary significantly depending on the design strategy employed.
  1. (1)
[…]
  2. (2)
    Beyond-Accuracy-Oriented Strategy. The goal of this beyond-accuracy-oriented approach is to broaden the user’s viewing experience by introducing diversity and novelty. It steers away from strictly aligning with known preferences and instead presents a variety of unique and perhaps surprising movie recommendations. It could be noted that for the same given query, this results in the user being recommended a movie like “Moon (2009)” - (genres: Drama—Mystery—Sci-Fi), which may be outside the user’s typical viewing history but offers a new cinematic perspective.
  3. (3)
    Reasoning-Oriented Strategy. This strategy is centered on enhancing user engagement and understanding. It not only recommends movies but also provides detailed explanation and reasoning regarding the suggestions. While using these explanations can potentially assist users in exploring and understanding why certain movies are recommended, we are particularly investigating this scenario since previous research has shown that motivating the LLM to reason, for example, through the use of a chain-of-thought (COT) reasoning, could result in more accurate responses from LLMs (i.e., relevant recommendations). An example of recommendation under this strategy could be “The Shawshank Redemption (1994)”, accompanied by an explanation of its themes of hope and redemption that resonate with the motivational depth found in “Toy Story.”
Promoting fairness in RecLLM: To combat popularity bias, we tested two approaches: (1) Instructional prompts: Statements like “Focus on fair recommendations, balancing popular and lesser-known movies” were integrated into various scenarios, aiming to influence RecLLM’s output. (2) Role-based directives: We explored directly incorporating fairness into the “system’s role”, testing its effectiveness compared to solely instructional prompts. This experiment seeks to identify the most effective strategies for achieving fair recommendations in RecLLM. In sum, our work demonstrates how strategically designing a prompt’s semantic intent and structure can impact a recommender system’s performance across accuracy, provider fairness, diversity, and temporal freshness (recency).
Contributions. The following is a list of proposed contributions in the current work:
  1. (1)
    Enhanced Analysis of Prompt Design in Zero-Shot RecLMMs. This research provides an in-depth study into how the design of prompts (in terms of intent, structure, and system role) affects the performance of ChatGPT-based recommendation systems. It investigates how different prompt design impacts factors such as accuracy, fairness, diversity, and temporal freshness in recommendations. This study involves experimental designs that compare various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and analyze their impacts on the recommendation of niche versus popular content, as well as the balance between old and recent movies.
  2. (2)
    Identification and Analysis of Different Biases in Zero-Shot RecLLMs. This work identifies and analyzes numerous biases that may be present in RecLLMs such as ChatGPT, thus laying the groundwork for investigating bias mitigation strategies. Key biases identified under scrutiny include popularity bias, genre preference bias, and temporal (recency) bias. We study how these biases uniquely manifest in LLM-based systems versus traditional Collaborative Filtering (CF) models.
  3. (3)
    Stability Analysis in Zero-Shot RecLLMs: This world also studies the stability of recommendations provided by LLM-based systems over time. Stability in this context refers to the consistency of recommendations given similar input conditions at different times. Our study analyzes and discusses factors that might impact stability (besides the parameters controlled by LLMs), such as changes in underlying data, model updates, and prompt variations. It would also explore the implications of stability for user satisfaction and trust in the system in real-life deployments.
The structure of the paper is organized as follows. Section 2 presents the related work, focusing on the exploration of fairness in recommendation systems (RS) and pre-trained language models (LMs), as well as their applications in enhancing RS. Section 3 is dedicated to presenting a thorough evaluation of ChatGPT. This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1), the repeated experiments for stability analysis (Section 3.2), an understanding of the system’s role in ChatGPT-based RecLLM (Section 3.3), and the promotion of fairness (Section 3.4). Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.5). Following this, Section 4 outlines the experimental setup employed in our study. The results and key findings of these experiments are thoroughly discussed in Section 5. The paper concludes with Section 6, where we summarize our findings and outline potential avenues for future research.
2. Related work
In this section, we briefly review some related work on recommendation systems and LLM techniques.
2.1. Fairness in Recommender Systems
Fairness has emerged as a pivotal topic within the AI community in the last few years, arguably becoming the most scrutinized aspect across various branches of trustworthy AI, including fairness, security, privacy, and explainability. In RS, fairness has garnered significant attention due to the multi-stakeholder nature of these systems (Deldjoo et al., 2023; Ekstrand et al., 2019). Unfairness, even in its minimal form, can adversely impact various stakeholders, including consumers, producers, system designers, supply chains, and even the environment – the latter are often referred to as ‘side-stakeholder’.
The body of literature on fairness in RS is diverse, encompassing multiple perspectives and dimensions. Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), have introduced various taxonomies to categorize these diverse aspects into several orthogonal or partially-orthogonal dimensions. Key dimensions recognized in the literature include the main stakeholder in question (e.g., consumer vs. producer), the target benefits associated with each (such as effectiveness vs. item exposure), the granularity of sensitive groups for assessing fairness (individual vs. group level), and other dimensions including the core definition of fairness, temporal aspects, and more. These dimensions offer a comprehensive view of the fairness landscape in RS, as detailed in these surveys.
In the context of the present study, and with regard to positioning this work within the literature of fair Recommender Systems (fair-RS), we have introduced Table 1. This table is designed to categorize existing literature along two principal dimensions: the stakeholder in question (consumer vs. producer) and the nature of the core RS under scrutiny (traditional RS vs. those based on recent advancements in LLMs). We briefly review these dimensions in the following.
Core RS under scrutiny. We distinguish between ‘traditional’ RS and those enhanced by ‘RecLLM’. This distinction is crucial because, while RecLLM, such as those based on GPT-like architectures, promise to significantly advance RS landscape with more nuanced and personalized recommendations, they also raise concerns about inherent biases. The vast and unregulated nature of internet data used for training LLMs raises concerns about inherent biases against specific races, genders, popular brands, and other sensitive attributes. For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands. Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways. Therefore, accurately measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies, ensuring that advancements in RecLLM do not amplify existing inequalities. This is the key goal of the current work in hand.
  • Traditional RS. This column lists studies that have focused on traditional methods of recommender systems. They primarily rely on CF algorithms (possibly using side information of users and items) without the advanced natural language processing capabilities of LLMs. Within these works, some are dedicated to building evaluation frameworks for evaluating RS unfairness, while others focus on developing various mitigation strategies.
  • RecLLM: This column spotlights the burgeoning research domain that merges language models (LMs) with RS representing a significant shift towards harnessing advanced NLP techniques to enhance the accuracy and relevance of recommendations. These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023a; Zhang et al., 2023). Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023a; Zhang et al., 2023), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
Table 1. Research landscape in Recommender Systems focusing on Consumer and Producer fairness with respect to Age, Gender, Other, and Others
Traditional RS RecLLM
Consumer Fairness Activity (Li et al., 2021; Hao et al., 2021; Xiao et al., 2020; Rahmani et al., 2022b)
Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023)
Merits (Sühr et al., 2021; Gómez et al., 2021)
Others (Wan et al., 2020; Lin et al., 2021)
Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023a)
Demographics (Kırnap et al., 2021; Boratto et al., 2021)
Price/Brand/Location (Deldjoo et al., 2021a; Burke et al., 2018; Liu et al., 2020; Shakespeare et al., 2020)
CP Fairness Mixed attributes (Naghiaei et al., 2022; Rahmani et al., 2022a; Chakraborty et al., 2017; Patro et al., 2020; Wu et al., 2021a; Do et al., 2021)
Stakeholder. As mentioned earlier, a dominant aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides. Within each of these segments, as you can observe, we can further categorize the literature based on which sensitive attribute groups are defined. For example, sensitive attributes such as the demographics of consumers (age, gender) and producers, as well as the popularity of items (on the produce side) are quite common focal points.
  • Consumer Fairness. This category is subdivided into various attributes like Activity, Demographics, Merits, and Others. It includes studies that focus on ensuring fairness among consumers of the recommender system based on these attributes. For instance, ensuring that recommendations are not biased towards a particular demographic group.
  • Producer Fairness. This focuses on the fairness towards the providers or producers of the content or products recommended by the system. It includes subcategories like Popularity, Demographics, and Price/Brand/Location. These studies might address issues like ensuring lesser-known or niche producers get fair visibility and opportunity in the recommendation process.
  • CP Fairness. This category involves studies that consider both consumer and producer fairness simultaneously, addressing the balance between the two.
It is noteworthy that while traditional Fair-RS research has been extensively explored, RecLLM is an emerging field, presenting its own unique considerations and challenges. Our work is situated in the ‘Producer Fairness’ category under ‘RecLLM’, focusing on producer fairness in the context of ChatGPT. It focuses on producer fairness in the context of ChatGPT, particularly examining how prompt engineering techniques can be leveraged to address or potentially enhance fairness.
The study by Zhang et al. (2023) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness. Their work introduces a novel benchmark, FaiRLLM, for evaluating the fairness of RecLLM, highlighting ChatGPT’s biases towards certain sensitive user attributes in music and movie recommendations. Conversely, the work by Li et al. (2023a) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation. This study investigates ChatGPT’s performance in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection. Similarly, research such as that by Di Palma et al. (2023) examines the personalization and popularity bias in RecLLMs, particularly in the context of ChatGPT and other LLMs. However, these studies do not address the intricacies of prompt engineering in RecLLMs, nor do they comprehensively address the various forms of biases studies in the current study. Our research goes beyond examining provider fairness based on popularity bias and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.
2.2. Leveraging Pre-trained LMs and Prompting for Recommender Systems
Recent advancements in recommender systems (RS) have been significantly influenced by the integration of LLMs and innovative prompting strategies. The use of natural language in recommendation tasks has been explored in various ways. For instance, Hou et al. (2022) utilize natural language descriptions and tags as inputs into LLMs to create user representations for more effective recommendations. This contrasts with the narrative-driven recommendations (Bogers and Koolen, 2017) that rely on verbose descriptions of specific contextual needs.
In terms of prompting techniques, early methods relied on few-shot prompting, where training examples are used as a guide for LLMs (Brown et al., 2020). With prompt learning, tasks are adapted to LLMs rather than the other way around, utilizing discrete prompts or continuous/soft prompts for task performance. This approach has shown promise across various tasks, including recommendation tasks.
Personalizing LLMs for recommendation is crucial for understanding a user’s intent and addressing their personalized needs. Recent efforts like P5 (Geng et al., 2022) and OpenP5 (Xu et al., 2023) have integrated several recommendation tasks into one LLM using personalized prompts. This approach reformulates recommendation tasks as sequence-to-sequence generation problems, demonstrating the flexibility of LLMs in handling diverse recommendation scenarios. Lastly, prompt transfer research, such as SPoT (Vu et al., 2021) and ATTEMPT (Asai et al., 2022), focuses on learning from source tasks and utilizing this knowledge for target tasks. This method, including knowledge distillation techniques, shows potential in intra-task prompt distillation and cross-task prompt transfer, contributing to the efficiency and effectiveness of LLM-based recommendation models.
3. Evaluation of ChatGPT-based RecLLM
Our work focuses on evaluating ChatGPT as a zero-shot recommender in the movie domain, aiming to extend beyond previous research by delving more deeply into biases and potential harms caused by such systems. Given the extensive nature of our experiments, we chose to concentrate on the single domain of movie recommendations. In this context, our evaluation methodology for ChatGPT was comprehensive and multifaceted. We conducted extensive experiments across a multi-dimensional testbed to assess the robustness and adaptability of the system. This evaluation particularly emphasized understanding the role and impact of ChatGPT in three key areas:
  1. (1)
    Goal-oriented prompt,
  2. (2)
    System role,
  3. (3)
    Promoting fairness, as well as
  4. (4)
    Implcit vs. explicit scenario
These dimensions were varied to observe their individual and combined effects on the quality of the recommendations provided. Please note that since the primary objective of this study is centered on (provider) fairness, the emphasis on fairness has been integrated either as a distinct scenario within the goal-oriented prompts or within the system role itself, and the impact is evaluated.
3.1. Goal-oriented prompts
Seven distinct prompt scenarios were designed, categorized into three classes focused on personalization, beyond-accuracy metrics, and reasoning. These classes aim to explore how variations in the nature of the prompt influence the recommendations. The scenarios are detailed in Table 2 and further exemplified in Section 1.
Table 2. Overview of Prompt Scenarios.
Scenario Description and Prompt
Personalization Focused
S1 - Simple
Basic recommendations without additional context.
Prompt: ”Recommend 10 movies that the user will likely enjoy.”
S2 - Genre-focused
Recommendations focusing on genres and themes similar to the user’s past favorites.
Prompt: ”Recommend 10 movies that the user will likely enjoy, particularly focusing on genres and themes similar to their past favorites.”
S3 - Rating-focused
Incorporation of the user’s explicit ratings into the recommendations.
Prompt: ”Recommend 10 movies the user will likely enjoy, taking into account both their favorite genres and past movie ratings.”
Beyond-Accuracy Focused
S4 - Diversify Recommendations
Suggesting lesser-known films to diversify the user’s experience.
Prompt: ”Suggest 10 high-quality, lesser-known films that diverge from mainstream blockbusters, yet align with the user’s tastes.”
S5 - Surprise
Offering unexpected recommendations for exploring new preferences.
Prompt: ”Offer 10 unique and unexpected movie recommendations aimed at broadening the user’s cinematic horizons beyond their usual preferences.”
Reasoning-Explanation Focused
S6 - Motivate Reasoning
Providing reasoning for each recommendation to enhance transparency.
Prompt: ”Provide 10 carefully selected movie recommendations, each accompanied by a rationale explaining its suitability for the user’s preferences.”
S7 - Chain of Thought (COT)
Engaging in a logical reasoning process to arrive at the recommendations.
Prompt: ”Let’s think this through: What would be 10 great movie recommendations for this user and why?”
In particular, scenarios S1 to S3 are focused on core personalization aspects, encompassing basic recommendations without additional context (S1), genre preferences (S2), and the incorporation of explicit user ratings (S3). Meanwhile, Scenarios (S4) and (S5) are designed to enhance the diversity and novelty of the recommendations, with S4 emphasizing lesser-known film recommendation and S5 aiming to broaden the user’s cinematic horizons with unique and unexpected choices. Scenarios S6 and S7 are explanation-motivation oriented, where S6 seeks to provide reasoning for each recommendation to enhance transparency, and S7 involves a logical ste-by-step reasoning process to arrive at the recommendations. These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022; Wang et al., 2022; Chu et al., 2023).
Note. The CF baselines used in this work operate in an implicit setting. For consistency and fairness, all scenarios in the generative part adhered to this setting, thereby not revealing user movie ratings to the LLM in the prompts. Scenario ‘Rating-focused’ S3 is the only exception, included solely for completeness. This is discussed in Section 3.5.
3.2. Repeated Experiment for the Stability of the Analysis
The experiments were conducted at a temperature setting of 0.0, aiming for predictable responses. However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in GPT’s behavior. This suggests that the recommendation system may interpret the same input differently in each iteration, independent of the temperature setting.
To assess the robustness and reliability of our system, we repeatedly conducted recommendation queries for each user, five times each, to examine if outputs vary despite consistent input.
We further adopted a “bootstrapping sampling strategy” during the evaluation stage, as detailed in Section 4.3. This method is particularly effective in revealing the reliability of the results, even when dealing with datasets that are small or may not perfectly represent the entire population. This thorough evaluation approach, therefore, enhances our confidence in the system’s performance across various potential real-world scenarios.
3.3. Understanding ChatGPT’s “System” Role
In this study, we sought to understand the impact of different system roles assigned to ChatGPT on recommendation outcomes. We focused on assessing whether responses significantly differ when specific “system” instructions are attached, and how these variations influence the functionality of a recommender system. Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system’s operations versus incorporating it in the prompt structure. Table 3 summarizes the roles assigned to ChatGPT during our experiments:
Table 3. Summary of Different Roles Assigned to ChatGPT during experiments
Role ID Description
R0 - No Role
Direct user-driven prompts without system context.
R1 - System Role as Recommender
System-centric role focusing purely on user information.
R2 - System Role as Fair Recommender
System role with an explicit fairness objective.
The roles explored provide insight into how ChatGPT’s framing, in terms of its responsibilities and objectives, shapes its behavior in making recommendations. This analysis is crucial for comprehending the subtleties of human-AI interaction within diverse contexts and for distinct purposes, especially in how the system interprets and adheres to varying directives.
3.4. Promoting Fairness
We also aim to understand whether integrating a fairness statement in the system or in the prompt itself is more effective. To assess the impact of explicit directives on the recommendation system, each scenario was optionally combined with a ‘Fairness Emphasis Statement.’ These statements directed the model towards specific objectives, here ensuring fairness. The considered emphasis options included
  • E0 - Without Fairness Statement: No fairness statement is included in this option. Prompts are used as described in Table 2.
  • E1 - With Fairness Statement: This option emphasizes ensuring fair recommendation between popuar nd less popular movies. “Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy”. The latter prompt is combining Scenario (S1/E1).
3.5. Explicit vs. Implicit Scenario
This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the recommender system’s quality, based on the premise that including such information could improve recommendations. Scenarios S1 to S6, except S3, were conducted implicitly, without user ratings in the prompts, mirroring the implicit mode of the CF baseline experiments. Uniquely, S3 was examined in both explicit and implicit contexts. In the explicit context, user ratings were integrated into the prompts, but were absent in the implicit context. This explicit approach was specific to S3. An example of {user_movies_string} and the presentation of user movie lists and ratings in the explicit context is as follows.
Example. To illustrate the experimental scenarios, consider a user’s movie list represented in this manner.
The Matrix (Genres: Action|Sci-Fi), Inception (Genres: Action|Thriller)
In the Explicit scenario, the prompt would be constructed to include these ratings, exemplified here:
The Matrix (Genres: Action/Sci-Fi, Rating: 5/5), Inception (Genres: Action/Thriller, Rating: 4.5/5)
The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendation more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
4. Experimental Setup
This section outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
4.1. Datasets and splitting
In this study, we utilized the MovieLens-Latest-Small dataset provided by the GroupLens group, [^1] https://grouplens.org/datasets/movielens/ which contains the most recent versions of movies. It is important to note that we intentionally focused our research within the domain of movies. This decision was made due to the potential varying capabilities of GPT-based models across different domains. Additionally, conducting extensive experiments with GPT models is time and resource-intensive, which is why we opted to work with this specific dataset. We acknowledge that there is potential for future exploration in other domains, and we leave this avenue open for future research. Detailed dataset parameters, including $\frac{R}{U}$ (ratings per user) and $\frac{R}{I}$ (ratings per item), resulting in a higher density $\frac{R}{U\times I}$, are presented in Table 4 for reference. Our setting involves a train-validation-test split for the data with ratios 80%, 10%, and 10%, respectively.
4.2. Evaluation Metrics
Our study employs a diverse range of metrics to assess the performance of our recommendation system. These metrics are categorized into two primary classes: Accuracy, Beyond-accuracy and Provider Fairness.
Accuracy Metrics. These metrics measure the quality of the top-$k$ ranking in the recommendation list.
  • NDCG@k. A measure of ranking quality, particularly the placement of relevant items.
  • Recall@k. The proportion of relevant items that are successfully recommended.
High scores in these metrics indicate more accurate and relevant recommendations.
Beyond-Accuracy and Provider Fairness Metrics. Beyond conventional accuracy metrics, these metrics evaluate the diversity, and fairness of the recommendations w.r.t provider perspective. They assess how well the system introduces new and varied content and how equitably it treats different item providers.
  • Gini Index: Measures the inequality in the distribution of item recommendations. A lower Gini Index indicates a more equitable distribution among items. The Gini Index is beneficial for identifying monopolization of recommendations by a few items, but it may not always capture the full picture of item diversity.
  • HHI (Herfindahl-Hirschman Index): Assesses the concentration of recommendations among items. A lower HHI value suggests a healthier, more diverse recommendation landscape. This index is particularly useful for detecting dominance by a small number of items, though it might overlook the distribution nuances among less recommended items.
  • Entropy: Reflects the diversity and unpredictability of recommendations. Higher entropy values indicate greater diversity. Entropy is advantageous for understanding the overall unpredictability of recommendations but may not directly indicate the fairness of item exposure.
Each of these metrics has its own advantages and limitations, making them collectively essential for a well-rounded evaluation. They help in understanding not just the effectiveness of the recommendation system in terms of user satisfaction (accuracy) but also its robustness in promoting diverse and fair content distribution.
4.3. Bootstrapping Sampling Strategy
In our evaluation stage, we employed a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement. This method was chosen to enhance the robustness and reliability of our analysis, particularly in measuring the performance metrics of our recommendation system, such as NDCG, Recall, and Precision.
Through this approach, we computed means for each metric across all bootstrap samples, then calculated their averages and 95% confidence intervals. This strategy is implemented for understanding the variability and confidence in our metrics, ensuring our evaluation reflects the system’s likely real-world performance. It is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
4.4. CF Baselines.
The choice of models and the corresponding hyperparameter optimization process is outlined as follows.
Table 4. Statistics of the final datasets used in this work after $k$-core pre-processing.
Dataset —U— —I— —R— $\frac{R}{U}$ $\frac{R}{I}$ Density Item Gini User Gini
MovieLens Latest Small 610 9,724 100,836 165.30 10.36 98.3% 0.715 0.603
4.4.1. Models.
We use a suite of competitive, CF recommendation models as baseline ranking models in our post-processing approach, as summarized below.
  • BPR (Rendle et al., 2012): A conventional recommendation model that employs matrix factorization to learn user and item embeddings of low dimensionality, and optimize the model based on the pairwise ranking of items for each user to predict whether a user prefers a given item over another;
  • ItemKNN (Sarwar et al., 2000): An item-based K-nearest employs neighbors algorithm, utilizing similarity metrics such as cosine similarity to identify the closest item neighbors. These neighboring items are then leveraged to predict scores for user-item pairs.
  • MultiVAE (Liang et al., 2018): A non-linear probabilistic deep learning model that extends a variational autoencoder (VAE) structure to collaborative filtering for implicit feedback, and acquires the underlying representations of users and items from their interactions to create recommendations in an unsupervised way.
  • LightGCN (He et al., 2020): A pure collaborative filtering method that utilizes a simplified version of graph convolutional networks (GCNs) without nonlinear activation functions and additional weight matrices. It learns user and item embeddings through graph propagation rules and user-item interactions, making it scalable and efficient.
  • NGCF (Wang et al., 2019): A graph-based recommendation model that employs a neural network architecture and learns high-order connectivity and user-item signals based on the exploitation of the user-item graph structure, by propagating embeddings on it.
The selected CF models constitute a set of competitive baselines from various natures, representing a diverse array of recommendation approaches (classical, neural, graph-based). Additionally, the TopPop method is included as a non-informative baseline, as an essential baseline to provide a case where the outcomes are heavily biased towards popular items.
4.4.2. Hyperparameter Tuning.
In our research, the primary focus is on the hyperparameter tuning of various collaborative filtering (CF) recommendation models, selected as baselines. These models are specifically tailored and evaluated within implicit feedback scenarios, where the subtleties of user preferences are inferred from indirect interactions. This approach ensures a comprehensive exploration of hyperparameter spaces to optimize model performance in these nuanced environments. The RecBole public library [^2] https://www.recbole.io/ is used for implementing and applying these models. Hyperparameter tuning is an essential step in optimizing the performance of these models. We employ a greedy search strategy to identify the best configurations based on their performance on the validation set. The following is a summary of the hyperparameter tuning conducted for each model:
  • BPR-MF. Embedding size choices ’[32, 64, 128]’ and learning rate choices ’[1e-4, 5e-4, 1e-3, 5e-3]’. Total of 12 cases.
  • ItemKNN. k choices ’[10, 50, 100, 200, 250, 300, 400]’ and shrink choices ’[0.0, 0.1, 0.5, 1, 2]’. Total of 35 cases.
  • NGCF. Learning rate choices ’[1e-4, 5e-4, 1e-3]’, hidden size list choices ’["[64, 64, 64]", "[128, 128, 128]"]’, node dropout choices ’[0.0, 0.1, 0.2]’, message dropout choices ’[0.0, 0.1, 0.2]’, and regularization weight choices ’[1e-5, 1e-3]’. Total of 108 cases.
  • MultiVAE. Learning rate choices ’[1e-4, 5e-4, 1e-3, 5e-3]’ and latent dimension choices ’[64, 128, 200, 300, 512]’. Total of 20 cases.
  • LightGCN. Learning rate choices ’[5e-4, 1e-3, 2e-3]’, number of layers choices ’[1, 2, 3, 4]’, and regularization weight choices ’[1e-05, 1e-04, 1e-03, 1e-02]’. Total of 48 cases.
The total number of hyper-parameters is 14, encompassing a total of 223 experimental cases. This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our resear
5. Results
In this section, we present insights gained from the experimental validation of our system. We have divided our analysis into four key areas, each focusing on different aspects of performance: (i) personalization (cf. Section 5.1), (iii) stability of recommendations over different runs (cf. Section 5.2), (ii) provider fairness and diversity of recommendations (cf. Section 5.3), and finally (iv) the temporal freshness (or recency) of recommendations (cf. Section 5.4). We also provide an interesting discussion on the genre dominance/bias of recommendation models employed in Section 5.5. We approach this section by formulating a number of the hypotheses for each part, afterward showcasing the results, and providing the associated analysis. Finally, we will provide answers to these hypotheses.
5.1. Personalization of Recommendations
This section focuses on the personalization performance measured by recommendation top-$K$ accuracy. Results can be found in Table 5, in which each generative scenario (i.e., using ChatGPT) could be found on the upper tabs while the result of CF baselines is on the bottom. It should be noted that, given the extensive nature of these experiments, not every possible scenario combination was tested independently; rather, a select subset of scenarios was examined.
We will analyze the results by evaluating the average performance (abbreviated as Avg Perf.) in terms of NDCG and Recall across different conditions/models. These conditions include exploring the impact of (i) the system’s role: No Role (R0), Standard Recommender (R1), and Fairness-aware Recommender (R2), and (ii) the presence (E1) or absence (E0) of a fairness statement in the LLM prompts.
Table 5. Table of the performance result measured in terms of recommendation Top-10 accuracy.
“System” Role: No (R0), Emphasis: No (E0)
Avg Perf. Bootstrap Mean Bootstrap Conf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.008803 0.00994 0.008784 0.010011 (0.006323, 0.011484) (0.00674, 0.013824)
S2. Genre-focused 0.006964 0.007672 0.006994 0.007689 (0.004881, 0.009200) (0.005079, 0.0106819)
S4. Diversify 0.001906 0.002946 0.001896 0.002946 (0.000822, 0.003185) (0.001559, 0.004582)
S5. Surprise 0.007265 0.006968 0.007286 0.00696 (0.004671, 0.010279) (0.004130, 0.010021)
S6. Motivate Reasoning 0.010152 0.010906 0.010073 0.01098 (0.007446, 0.013124) (0.007511, 0.0148920)
S7. Chain-of-thought (COT) 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
“System” Role: No (R0), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
Type Model NDCG Recall NDCG Recall NDCG Recall
Generative (ChatGPT) S1. Simple 0.010431 0.010557 0.010404 0.010579 (0.007491, 0.013792) (0.007295, 0.014063)
S2. Genre-focused 0.005221 0.005471 0.005211 0.00555 (0.003413, 0.007300) (0.003361, 0.008196)
S4. Diversify 0.001717 0.003029 0.001713 0.003024 (0.000607, 0.003261) (0.001504, 0.005083)
S5. Surprise 0.004095 0.005089 0.004059 0.005112 (0.002399, 0.006132) (0.002645, 0.008248)
S6. Motivate Reasoning 0.004315 0.005097 0.004302 0.005031 (0.002246, 0.006819) (0.002668, 0.007909)
S7. Chain-of-thought (COT) 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
“System” Role: Normal Recommender (R1), Emphasis: Fair (E1)
Avg Perf. Avg. Bootstrap Bootstrap Conf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.013007 0.012606 0.012984 0.012514 (0.009799, 0.016379) (0.008610, 0.016971)
S2. Genre-focused 0.00854 0.00987 0.008582 0.00983 (0.006277, 0.010984) (0.006564, 0.013434)
S4. Diversify 0.00314 0.005606 0.003162 0.005523 (0.001477, 0.005548) (0.003086, 0.008230)
S5. Surprise 0.006473 0.006435 0.006403 0.00642 (0.004065, 0.009319) (0.003627, 0.009950)
S6. Motivate Reasoning 0.009379 0.00957 0.009447 0.009496 (0.006160, 0.012722) (0.006056, 0.013061)
S7. Chain-of-thought (COT) 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
“System” Role: Fair Recommender (R2), Emphasis: No (E0)
Avg Perf. Avg. Bootstrap Bootstrap Cnf
Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
Generative (ChatGPT) S1. Simple 0.014268 0.015796 0.014257 0.015731 (0.011243, 0.017614) (0.011896, 0.019893)
S2. Genre-focused 0.008847 0.01132 0.00883 0.011382 (0.006700, 0.011233) (0.007998, 0.015372)
S4. Diversify 0.002844 0.004901 0.002825 0.004872 (0.001303, 0.004922) (0.003205, 0.006746)
S5. Surprise 0.007504 0.00729 0.007446 0.007304 (0.004587, 0.010622) (0.004066, 0.011234)
S6. Motivate Reasoning 0.012354 0.011968 0.012407 0.011929 (0.008901, 0.016671) (0.008133, 0.016052)
S7. Chain-of-thought (COT) 0.013212 0.012862 0.013162 0.012908 (0.009972, 0.017185) (0.009256, 0.016893)
CF Baselines
Discriminative BPR-MF 0.047762 0.060699 0.047742 0.060537 (0.040465, 0.054958) (0.050181, 0.071453)
ItemKNN 0.04905 0.065558 0.048995 0.065367 (0.041122, 0.056398) (0.054682, 0.076526)
NGCF 0.04445 0.053534 0.044376 0.053448 (0.037306, 0.051784) (0.044374, 0.063413)
VAE 0.047158 0.056589 0.047187 0.056502 (0.039967, 0.054677) (0.047093, 0.065803)
LightGCN 0.048295 0.059108 0.048215 0.059019 (0.041621, 0.055491) (0.049809, 0.068851)
TopPop 0.030556 0.032327 0.030539 0.032127 (0.024253, 0.036734) (0.025213, 0.039704)
Additionally, a comparison between the regular performance values (obtained using average across values) and bootstrap averages of model performances provides deeper insights. The bootstrap method, which involves repeatedly resampling with replacement, offers a more robust understanding of the variability in performance measures. Table 5 illustrates this comparison. For instance, while the regular average provides a single estimate of NDCG and Recall, the bootstrap mean extends this by offering confidence intervals. These confidence intervals, represented as Bootstrap Conf. in the table, indicate the range within which the true performance measure is likely to fall with a certain level of confidence.
We outline below the hypotheses regarding personalization that are central to this section:
Hypothesis 1 (H1): The incorporation of a variety of goal-oriented prompts (as detailed in scenarios S1 to S7 in Section 3.1) across diverse settings (e.g., involving different system assignment roles) is anticipated to yield substantially varied performance outcomes in RecLLM. Hypothesis 2 (H2): GPT-based recommenders, particularly in a zero-shot setting, could achieve results comparable to CF baselines in specific scenarios. This is more probable in situations where the system is encouraged to engage in motivated reasoning and/or adopt COT strategies. This hypothesis is founded on observations that such strategies have demonstrated competitive performance in previous ML tasks.
The Impact of Prompt Structure and Setting
In the uppermost tab of Table 5, which presents the settings designated as No Role (R0) for the system role, the results are typically less favorable. Specifically, the prompt scenarios Simple (S1) and Motivated Reasoning (S5) demonstrate better performance within their respective segments. Scenario S1 under No Role (R0) reports NDCG and Recall values of 0.008803 and 0.00994, respectively. In comparison, S5 shows slightly improved metrics, with an NDCG of 0.010152 and a Recall of 0.010906. Overall, these results are the poorest when compared to other settings utilizing LLM-promoted prompts.
Contrastingly, the data in tabs 3 and 4 indicate that when roles are delineated as either Normal Recommender (R1) or Fair Recommender (R2), there is a notable enhancement in both NDCG and Recall metrics. It is of particular note that under these roles (R1 and R2), scenarios such as simple (S1), chain-of-thought (COT) (S7), and Motivated Reasoning (S6) markedly exceed the performance metrics of their No Role (R0) counterparts. The most pronounced performance increase is evident in the fourth tab, where models reach their optimal enhancements. The simple model (S1) advances from 0.008803 to 0.014268 in NDCG (approximately a 62% increase) and from 0.00994 to 0.015796 in Recall (approximately a 59% increase). The Motivated Reasoning (S6) and COT (S7) models also exhibit substantial gains, with S6 escalating from 0.010152 to 0.012354 in NDCG (21.7% increase) and from 0.010906 to 0.011968 in Recall (9.7% increase), and S7 rising from 0.006573 to 0.013212 in NDCG (101% increase) and from 0.005848 to 0.012862 in Recall (120% increase). This trend is similarly observed in the Normal Recommender, Emphasis: Fair (R2/E1) configuration. These results emphasize the significant advantages of integrating system roles into recommendation systems.
Bootstrap values in the R2-E0 setup further affirm the consistency and reliability of these improvements. Both the ‘simple’ and ‘COT’ models exhibit stable performance ranges across different roles and conditions.
Comparison with CF Baselines
When evaluating the effectiveness of GPT-based recommendation models compared to CF baselines, it is apparent that both the NDCG and the recall values for the GPT models are generally lower than those for the CF baselines. This suggests that, while GPT-based models have made significant strides, especially when designated with specific system roles such as “Normal Recommender” (R1) or “Fair Recommender” (R2), they still lag behind CF models in terms of both ranking relevant items at the top (NDCG) and retrieving a broader set of relevant items.
For example, even with the “Fair Recommender” (R2) role applied, the “Simple” (S1) GPT-based model achieves an NDCG of 0.014268 and a recall of 0.015796, which are enhancements of the scenario “No Role” (R0) but still do not surpass the CF baseline models such as “BPR-MF,” “ItemKNN,” “NCF,” and “LightGCN,” that demonstrate higher efficacy in both metrics. The “TopPop” model, while having the lowest scores among the CF baselines, still presents an NDCG and Recall that are competitive with the improved GPT-based models.
In summary, some noteworthy best and worst performances can be found below:
  • For the GPT-based models under the “Fair Recommender” (R2) role with no fairness emphasis (E0):
    • The Simple (S1) scenario achieves the best NDCG of 0.014268 and a Recall of 0.015796.
    • The Diversify (S4) scenario shows the worst NDCG of 0.002844 and a Recall of 0.004901.
  • For the CF baselines:
    • The ItemKNN model has the highest performance with an NDCG of 0.04905 and a Recall of 0.06558.
    • The TopPop model displays the lowest performance among the CF baselines with an NDCG of 0.030556 and a Recall of 0.032327.
Despite the clear impact of prompts on the performance of RecLLM, it can be noted that Top-$k$ recommendation performance is not on par with traditional CF baselines. There are multiple strategies to address this issue, including enhancing prompt clarity and transitioning to few-shot learning. However, we posit that a significant factor contributing to this discrepancy is the possibility that many movies recommended by the GPT-recommender might not be present in the dataset. Therefore, by constraining GPT to select items only from the test item set, we believe that this could markedly improve the accuracy and relevance of the recommendations.
H1: Supported. The introduction of specific roles (e.g., Normal and Fair Recommender) in recommendation systems significantly enhances the performance of generative models. This is evidenced by substantial increases in NDCG and Recall metrics when comparing the No Role scenario (R0) with Role-Based scenarios (R1, R2). It is worth noting in terms of which prompt scenario is effective, the Simple (S1) scenario, suggesting “Recommend 10 movies that the user will likely enjoy,” appears to provide the most effective approach for achieving the best top-$K$ recommendation accuracy. In other scenarios, only the reasoning scenario, such as COT (S7) matches this performance. (H2): Partially Supported. The evidence does not support Hypothesis 2 (H2), which posited that GPT-based recommenders could match CF baselines in certain settings. While GPT models, particularly with system roles and strategies scuh as Simple, ‘Motivated Reasoning’ and ‘COT’, show enhanced NDCG, Recall, they generally do not reach the performance of CF baselines. It is worthwhile that these results indicate a trend towards achieving outcomes closer to those of Top-Pop with the refinement of prompt design, better rationalizing, or transitioning to a few-shot learning setting.
5.2. Stability and Confidence of Personalization Metrics
As stated in Section 3.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models. To study the impact of such randomness, each experiment was conducted five times specifically for GPT-based models. It is important to note that prior to our experiments, we lacked specific knowledge about the performance of Chat-GPT prompts. Therefore, our experiments focused exclusively on the R0/E0 scenario, which demonstrated poorer performance compared to the CF baseline. Despite this, the variability of results across different runs is noticeable. Consequently, in the analysis of this section, we should focus on observing the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
Hypotheses on Stability of the Top-$K$ Accuracy Metric
Hypothesis 1 (H1): Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent NDCG scores.
The graph in Figures 2 illustrates the NDCG and Recall scores for GPT-based recommendation models across five different runs. These runs overall have been conducted at different points in time that could span from hours to a few days. Notably, there is observable variability in the performance metrics from run to run. For instance,
  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std333standard deviation of approximately 0.000305.
  • Motivate reasoning (S6): The NDCG scores are [0.0101, 0.0097, 0.0093, 0.0094, 0.0095], with a std of approximately 0.000316.
  • COT (S7): The NDCG scores are [0.0066, 0.0067, 0.0070, 0.0069, 0.0068], with a std of approximately 0.000158.
Despite the inherent randomness in individual iterations, the general tendency of the results remains consistent, suggesting that while individual runs may produce fluctuating scores, the overall behavior of the models does not significantly deviate. This is evidenced by the low standard deviations in the NDCG scores (0.000305 for S1, 0.000316 for S6, and 0.000158 for S7), indicating a high level of result stability over different runs. Such consistency aligns with the expected behavior of generative models, which can exhibit some degree of randomness in their output due to the stochastic nature of the underlying data/algorithms. However, the close clustering of these scores, as reflected in the narrow stds, indicates a stable pattern of performance.
(a) NDCG@10.
(b) Recall@10
Figure 2. Studying the stability of GPT-based performance metric across different runs.
While the current study demonstrates a high degree of stability in model results over multiple runs, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. The dynamic nature of data and algorithms in generative models means that updates and changes over time could lead to different behaviors and outputs. Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the model’s performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
H1: Supported. The data presented in the NDCG graph aligns with the hypothesis. While there is some expected variability across multiple runs, it remains within an acceptable/predictable range. Confidence intervals overlap significantly for different runs of each GPT-based model. While this consistency across runs confirms the system’s reliability in providing personalized recommendations, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months. Therefore, it is advisable to approach these results with caution when considering long-term deployment.
5.3. Diversity and Provider Fairness
This section examines the diversity and fairness of different recommender models. Fairness is assessed by measuring the equitable distribution of recommendations across all items, avoiding bias towards popular or specific subsets (Gini Index, HHI). Diversity is gauged through the variety and uniqueness of the recommendations, ensuring the inclusion of a broad spectrum of items (Entropy), refer to Section 4.2. These evaluations are crucial to understanding each model’s effectiveness in matching user preferences and its capacity to foster a balanced and varied recommendation environment.
H1: Given their training on extensive internet data, GPT-based models are hypothesized to exhibit a greater popularity bias than CF models, potentially leading to lower diversity in recommendations. This will be assessed using the Gini Coefficient and HHI for fairness, and Entropy for diversity. H2: The inclusion of a Fair Recommender as a ”system” role in GPT-based systems is expected to mitigate popularity bias and enhance the diversity of the system’s output.
For H1, upon examining the results presented, in Table 6 a notable observation is that the ItemKNN model consistently exhibits the highest recommendation fairness (achieving the lowest Gini Coefficient and highest Entropy) among all models. This, to conduct a focused discussion, we focused on the remaining machine learning-based CF models and GPT-based models, we will exclude ItemKNN from subsequent comparisons.
Without considering ItemKNN, the general trend suggests that the top-performing GPT models (Simple, Motivate Reasoning, and Chain-of-thought (COT)) are somewhat comparable to the majority of ML-based CF models, except when compared to NGCF, which shows a notably better performance in terms of fairness (Gini Coefficient: 0.950845 vs. 0.0982346, Entropy: 6.420996 vs. 5.042821) and diversity. Intriguingly, methods designed to promote diversification and novelty in GPT case (Surprise and Diversify Recommendation) perform contrarily, evidenced by their increased Gini Coefficients (Surprise: 0.997906, Diversify Recommendation: 0.992349) and reduced Entropy (Surprise: 3.227737, Diversify Recommendation: 4.232139), which is almost the worst in the respective scenarios, suggesting that these strategies may not effectively align with the objectives of reducing bias and enhancing diversity.
Therefore, upon excluding the exceptional performance of ItemKNN, which seems to leverage its distinct mechanism effectively in a dataset characterized by popularity bias, the GPT-based models’ levels of fairness and diversity appear comparable to those of CF models. The data suggests that GPT-based models generally match their CF counterparts, with exceptions such as NGCF which shows superior fairness. Interestingly, NGCF, despite being the least performing CF model in some aspects, still outperforms GPT-based models on fairness. When we turn our attention to the top-performing CF models, namely BPR and LightGCN, it is evident that they achieve their enhanced performance by concentrating user preferences on a narrower selection of items, thus limiting diversity—a trend less likely to be observed in GPT models. This is attributable to the GPT models’ design, which is not inherently influenced by the popularity bias present in our datasets.
For H2, the adaptation of GPT-based models with a ‘Fair Recommender’ system role results in a significant improvement in fairness and diversity metrics. Excluding ItemKNN for its good performance, the top GPT models exhibit better fairness in the Fair Recommender scenario, as evidenced by a noticeable decrease in Gini Coefficient (Simple: from 0.982463 to 0.978925, Chain-of-thought (COT): from 0.986889 to 0.979313) and HHI, coupled with an increase in Entropy (Simple: from 5.042821 to 5.387465, Chain-of-thought (COT): from 4.619500 to 5.365294). This trend suggests that GPT-based models, when guided by the system role within the ChatGPT framework, demonstrate a promising capacity for controllability. Such adaptability could mitigate potential biases inherent in the models—a flexibility that may not be as pronounced in traditional CF models.
However, specific strategies such as Diversification (S4) and Surprise (S5) integrated into GPT-based models do not uniformly follow this positive trend. For example, in the Diversify scenario, we observe a slight increase in the model’s Gini coefficient (from 0.992349 to 0.992603) and a marginal rise in HHI (from 0.034724 to 0.030010), which deviates from the expected reduction. Similarly, the Surprise scenario shows a rise in Gini Coefficient (from 0.997906 to 0.998365) and HHI (from 0.059857 to 0.067952), alongside a reduction in Entropy (from 3.227737 to 3.023948). These anomalies highlight that certain approaches within the Fair Recommender role may inadvertently intensify the biases they aim to alleviate.
The broader implication of this study is that the system role in the ChatGPT framework, alongside certain prompt scenarios, has the potential to align or even surpass the fairness and diversity benchmarks set by classical CF models. In notable instances, the bias is reduced, as indicated by lower Gini and HHI values, and increased Entropy. This indicates a potentially significant step forward in the controllability of GPT-based models to prevent their biases, an attribute that arguably gives them an edge over CF models in managing the diversity and fairness of recommendations.
Answer to H1: Partially supported. The results do not uniformly support H1. While certain GPT-based models do exhibit higher popularity bias as indicated by the Gini Coefficient and HHI compared to some CF models, this is not a consistent finding across all models. For example, models such as NGCF show superior fairness indicators compared to some GPT-based models. However, when excluding the outlier performance of ItemKNN, GPT-based models are generally on par with CF models regarding fairness and diversity. Therefore, the hypothesis that GPT-based models inherently lead to lower diversity due to popularity bias is only partially supported. Answer to H2: Supported. The evidence supports H2 to a significant extent. When GPT-based models adopt a Fair Recommender role, there is a noticeable improvement in fairness and diversity metrics, such as a decrease in Gini Coefficient and HHI, along with an increase in Entropy. This indicates that GPT-based models have an inherent flexibility that, when leveraged with the right system role, can effectively mitigate popularity bias and enhance the diversity of recommendations. In both instances, it is critical to acknowledge that specific strategies, namely “diversification” and “surprise,” do not conform to their anticipated effectiveness—in some cases, they produce outcomes contrary to expectations. This observation implies that certain types of prompts may inadvertently introduce complexity, leading to a confounding of the system’s performance.
Table 6. Diversity and provider fairness of the experimented models
Model Normal Recommender Fair Recommender
Gini Coefficient HHI Entropy Gini Coefficient HHI Entropy
Simple 0.982463 0.017204 5.042821 0.978925 0.010899 5.387465
Genre-focused 0.964743 0.006455 5.919697 0.959879 0.004771 6.110040
Diversify Recommendation 0.992349 0.034724 4.232139 0.992603 0.030010 4.321307
Surprise 0.997906 0.059857 3.227737 0.998365 0.067952 3.023948
Motivate Reasoning 0.981745 0.019189 5.026322 0.979133 0.011218 5.366627
Chain-of-thought (COT) 0.986889 0.027030 4.619500 0.979313 0.011167 5.365294
BPR-MF 0.991758 0.012550 4.658056 0.991758 0.012550 4.658056
Item-KNN 0.914271 0.002877 6.671847 0.914271 0.002877 6.671847
NGCF 0.950845 0.002762 6.420996 0.950845 0.002762 6.420996
VAE 0.989722 0.009554 4.903511 0.989722 0.009554 4.903511
LightGCN 0.989610 0.010546 4.861879 0.989610 0.010546 4.861879
TopPop 0.994859 0.020000 3.912023 0.994859 0.020000 3.912023
5.4. Temporal Freshness Bias in Movie Recommendations
In this section, we study the recency and freshness of the movie recommendations provided by various GPT-based and CF models. The investigation focuses on whether the models demonstrate a bias toward recommending newer or older films. By examining the mean and median release years of the movies suggested by each model, we can gauge their tendency toward recency.
Figure 3. Distribution of Movie Release Years as recommended by different models
Table 7 presents the statistics that inform our understanding of each model’s tendency towards recommending newer or older films. Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007. This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense. Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself. In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films.
This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired. The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
Table 7. Model Statistics Reflecting Recency Bias
Model Median Year Std Year
Simple 1999 15.09
Genre-focused 1997 14.99
Diversify 2007 11.76
Surprise 2006 10.62
Motivate reasoning 2002 15.71
COT 1999 15.82
BPR-MF 1995 8.27
ItemKNN 1995 12.14
NGCF 1995 12.25
VAE 1995 8.50
LightGCN 1995 8.34
Pop 1995 5.65
Answer to H1: Contrary to the hypothesis, the results reveals that GPT-based models, especially ‘Diversify’ (Median Year: 2007) and ‘Surprise’ (Median Year: 2006), tend to recommend significantly newer movies than CF models. CF models such as BPR-MF (Median Year: 1995) and Pop (Median Year: 1995) generally suggest older films, as reflected in the lower mean and median years. The larger standard deviation in years for GPT-based models like ‘Surprise’ (Std Year: 10.62) also indicates a broader temporal range in their recommendations. This suggests that GPT-based models, perhaps due to their ability to capture more recent trends and discussions from diverse datasets, are capable of recommending more recent content than CF models, thus refuting the initial hypothesis.
5.5. Analysis of Genre Bias/Affinity in Recommender Systems
The goal of this section is to examine the genre biases present in movie recommendations provided by Collaborative Filtering (CF) models and GPT-based models. By comparing the prevalence of genre terms in the recommendations, we aim to understand if a particular model, or class of models, has a specific genre affinity (or bias) which could illuminate the performance differences between these models.
H1: We hypothesize that the differences in recommendation performance between CF and GPT-based models may be attributed to their respective genre preferences. CF models are expected to closely adhere to genre affinity in users’ profiles, while GPT models may exhibit a predominant recommendation of certain genres over others.
Upon analyzing the word clouds, it is evident that CF models such as BPR-MF, LightGCN, and VAE predominantly recommend genres such as “Action,” “Adventure,” “Thriller,” and “Sci-Fi.” These genres appear frequently and in larger font sizes, indicating their dominance in the recommendations made by these models.
In contrast, GPT-based models, represented in the lower word clouds, show a more varied genre distribution with “Comedy,” “Drama,” and “Romance” and even “Thriller” and “Adventure” being more prominent. These results are interesting and suggest that GPT-based RecLLMs may be less constrained by the users’ historical genre affinities and more exploratory in their recommendations, possibly leading to a broader but potentially less precise genre match to individual user profiles.
Figure 4. WordCloud of Movie Genres as recommended by different models. Top model correspond to CF models (BPR-MF, LightGCN, RecVAE), while lower models include GPT-based recommenders (Simple, Diversity, COT).)
Answer to H1: The analysis supports the hypothesis to a certain extent. While CF models tend to reinforce existing user preferences by frequently recommending popular genres such as “Action” and “Sci-Fi,” GPT-based models demonstrate a broader genre coverage in their recommendations. The presence of a more diverse set of genres like “Comedy,” “Drama,” and “Romance” in GPT-based recommendations suggests these models may be utilizing their extensive training on diverse data sources to introduce a wider variety of content. This could potentially lead to a discovery of new interests for users, though it may sometimes diverge from their established preferences. Such a tendency for broader genre representation by GPT-based models could be a contributing factor to their distinct performance characteristics compared to CF models.
5.6. Implicit vs. explicit scenario
Analyzing the impact of including ratings in the prompt shows noteworthy results. In our study, we initially avoided using ratings for the GPT models, aligning with the implicit setting of the Collaborative Filtering (CF) baseline. However, for completeness, we explored the effect of revealing user ratings for selected items.
The “Include-rating” model demonstrates a significant improvement compared to other GPT-based scenarios. With an NDCG of 0.013230 and a Recall of 0.014721, it surpasses genre-focused, Diversify Recommendation, Surprise, Motivate Reasoning, and chain-of-thought (COT) models. The Bootstrap NDCG and Recall for the Include-rating model are 0.013280 and 0.014673, respectively, indicating consistency in performance. The confidence intervals for NDCG (0.009737 to 0.017012) and Recall (0.010261 to 0.019795) further emphasize its reliability.
These results suggest that including ratings significantly enhances the recommendation system’s performance. It provides a more personalized and accurate reflection of user preferences, which is a key aspect in recommendation systems. While other GPT-based models showed varied performance, the inclusion of ratings seems to offer a more consistent and improved outcome. This insight could be valuable in refining other scenarios and models, potentially leading to better overall system performance.
6. Conclusion and Future Directions
This research presents a comprehensive analysis of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems. The study emphasizes the nuanced capabilities and inherent biases of these systems, highlighting their contrasting behaviors compared to traditional CF in the domain of movie recommendations. Our experimental findings underscore the significant impact of ‘prompt design’ strategies on recommendation quality, touching upon aspects such as accuracy, provider fairness, diversity, temporal stability, genre dominance, and freshness (recency).
We discovered that role-based prompts (those using clear system roles such as you a Recommender, or you are a Fair recommender) enhance fairness and diversity, effectively mitigating popularity bias. Interestingly, GPT-based models, while not always paralleling the performance of CF baselines, exhibit a preference for recommending newer and more diverse movie genres.
Moreover, our experiments reveal that including ratings in prompts appears to improve the personalization aspect of recommendations. GPT-based models also have a distinct inclination towards recommending newer, particularly those released post-2000. We hypothesize that a key reason for certain discrepancies in recommendations quality is that the GPT model might suggest movies not included in the dataset. By limiting the GPT model to choose only from a predefined set of test items, we anticipate a notable improvement in both the accuracy and relevance of its recommendations. Additionally, transitioning to a zero-shot scenario could significantly enhance personalization results in this context. In genral, we think that it important to balance personalization power of CF models with the broader exploratory nature of GPT-based RecLLMs, to ensure a diverse and engaging user experience.
Key areas for future research include:
  1. (1)
    Refine Prompt Design Strategies and Dataset Constraints. This involves developing prompts that effectively balance user preferences with broader content exploration, while also constraining the system to select items only from within the dataset. This limitation is expected to improve the accuracy and relevance of recommendations.
  2. (2)
    Explore Few-shot Learning to Improve accuracy and Contextual Relevance. Few-shot learning involves exposing the model to a small number of examples from the target domain, which can enhance its ability to make more informed recommendations.
  3. (3)
    Integrate GPT with CF in a RAG framework. Aimed to create a hybrid approach that leverages the strengths of both methodologies. This combination could potentially lead to more robust and effective recommender systems.
  4. (4)
    Expand research to cross-domain applications and conduct long-term user engagement studies. This includes evaluating RecLLMs in various content domains (e.g., books, music, e-commerce) and studying user interaction over extended periods to gauge satisfaction, trust, and the evolving nature of user preferences.
By focusing on these key areas, future research can significantly advance the field of recommender systems, particularly in leveraging the capabilities of large language models like GPT in diverse and dynamic recommendation scenarios.

### v3

@@
 Abstract.
-This study explores the nuanced capabilities and inherent biases of Recommender Systems using Large Language Models (RecLLMs), with a focus on ChatGPT-based systems.
-It studies into the contrasting behaviors of generative models and traditional collaborative filtering models in movie recommendations.
-The research primarily investigates prompt design strategies and their impact on various aspects of recommendation quality, including accuracy, provider fairness, diversity, stability, genre dominance, and temporal freshness (recency).
-Our experimental analysis reveals that the introduction of specific system roles and prompt strategies in RecLLMs significantly influences their performance.
-For instance, role-based prompts enhance fairness and diversity in recommendations, mitigating popularity bias.
-We find that while GPT-based models do not always match the performance of CF baselines, they exhibit a unique tendency to recommend newer and more diverse movie genres.
-Notably, GPT-based models tend to recommend more recent films, particularly those released post-2000, and show a preference for genres like ‘Drama’ and ‘Comedy,’ and ‘Romance’ (compared to CF ‘Action’, ‘Adventure’) presumably due to the RecLLMs’ training on varied data sets, which allows them to capture recent trends and discussions more effectively than CF models.
-Interestingly, our results demonstrate that the Simple and Chain of Thought (COT) paradigms yield the highest accuracy.
-These findings imply the potential of combining these strategies with scenarios that favor more recent content, thereby offering a more balanced and up-to-date recommendation experience.
-This study contributes significantly to the understanding of emerging RecLLMs, particularly in the context of harms and biases within these systems.
-It highlights the complexities and potential of prompt design in impacting recommendation outcomes and paves the way for future research in optimizing accuracy, fairness, and user experience in such systems.
+This paper explores the biases inherent in ChatGPT-based recommender systems, focusing on provider fairness (item-side fairness).
+Through extensive experiments and over a thousand API calls, we investigate the impact of prompt design strategies—including structure, system role, and intent—on evaluation metrics such as provider fairness, catalog coverage, temporal stability, and recency.
+The first experiment examines these strategies in classical top-K recommendations, while the second evaluates sequential in-context learning (ICL).
+In the first experiment, we assess seven distinct prompt scenarios on top-K recommendation accuracy and fairness.
+Accuracy-oriented prompts, like Simple and Chain-of-Thought (COT), outperform diversification prompts, which, despite enhancing temporal freshness, reduce accuracy by up to 50%.
+Embedding fairness into system roles, such as “act as a fair recommender,” proved more effective than fairness directives within prompts.
+We also found that diversification prompts led to recommending newer movies, offering broader genre distribution compared to traditional collaborative filtering (CF) models.
+The system showed high consistency across multiple runs.
+The second experiment explores sequential ICL, comparing zero-shot and few-shot learning scenarios.
+Results indicate that including user demographic information in prompts affects model biases and stereotypes.
+However, ICL did not consistently improve item fairness and catalog coverage over zero-shot learning.
+Zero-shot learning achieved higher NDCG and coverage, while ICL-2 showed slight improvements in hit rate (HR) when age-group context was included.
+Overall, our study provides insights into biases of RecLLMs, particularly in provider fairness and catalog coverage.
+By examining prompt design, learning strategies, and system roles, we highlight the potential and challenges of integrating large language models into recommendation systems, paving the way for future research.
+Further details can be found at https://github.com/yasdel/Benchmark_RecLLM_Fairness.
 Recommender Systems, Large Language Models, Bias and Fairness in RS, Movie Recommendation Analysis, Prompt Design Strategies, ChatGPT, Stability and Diversity in Recommendations
@@
 1. Introduction
-Recommender systems have increasingly become integral in various domains, offering substantial benefits for stakeholders in two-sided, or broadly, multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019).
-These systems alleviate user information overload and enhance producer sales and revenue.
-From a learning viewpoint, recommendation models can generally be classified into two categories: ‘discriminative’ and ‘generative’ models (Deldjoo et al., 2022).
-Discriminative models, often considered the traditional approach in recommendation systems, function by categorizing and differentiating to inform decisions about user preferences.
-For example, in the movie domain, these models can be compared to expert critics that are adept at discerning the distinctive characteristics that separate one type of item (movie) from another, given the audience.
-In other words, this differentiation is informed by factors such as user historical ratings, genre preferences, and other relevant attribute of items.
-Mainstream collaborative filtering (CF) recommendation models, such as matrix factorization (MF)-based models and more complex neural variations such as NeuMF (He et al., 2017), operate on this principle.
-Similarly, graph-based models, such as NGCF (Wang et al., 2019) and LightGCN (He et al., 2020), leverage the relational structure of the data to improve the quality of recommendations.
-Regardless of the methodological approach, traditional discriminative models are heavily based on knowledge derived from limited datasets and specific domains to calculate recommendations and inform decisions.
-Generative models represent a transformative approach, focusing on understanding the underlying structure and distribution of data to generate authentic-looking content.
-Their application in recommender systems (RS) exemplifies a shift from simple preference matching to a more nuanced understanding of user behavior and interests.
-In the context of movie recommender systems, these models can be likened to hypothetical curators who not only distinguish between various items but also comprehend the overarching patterns and tastes of the audience.
-Thus, these models go beyond mere differentiation; they synthesize and anticipate user preferences, providing tailored recommendations.
-Large Language Models (LLMs), exemplified by the Generative Pretrained Transformers (GPT), recently heralded a new era in Natural Language Processing (NLP).
-These Transformer-based models, such as BERT, GPT series, PaLM series, etc., are characterized by their vast number of parameters and are trained on extensive datasets using self/semi-supervised learning techniques.
-Their training on a wide range of internet text data endows LLMs with the exceptional ability to generate contextually relevant and coherent text in response to input prompts.
-Different from traditional recommendation systems, LLM-based models excel in capturing contextual information, comprehending user queries, item descriptions, and other textual data more effectively (Geng et al., 2022).
-Consequently, specifically-designed LLMs for recommender systems (RecLLMs) can harness LLMs to extract high-quality representations of textual features and leverage extensive external knowledge encoded within LLMs (Liu et al., 2023) to offer more nuanced insights into understanding user preference and profile.
-This is a significant leap forward compared to traditional RS.
-This aligns well with the overarching goal of delivering personalized, context-aware recommendations that reflect individual tastes and behaviors.
-In general, LLMs can contribute to recommender systems in three principal ways (Wu et al., 2023; Li et al., 2023b; Xu et al., 2024).
-First, LLMs serve as direct recommenders by generating personalized content and suggestions, stemming from their advanced understanding of context and user queries (Geng et al., 2022; Wu et al., 2023).
-Second, they enhance traditional models by extracting rich semantic representations from textual data.
-This aspect is particularly important in addressing the data sparsity issue in historical interactions (Da’u and Salim, 2020).
-Third, LLMs act as complex simulators in the recommendation process, effectively testing and refining the system’s predictions (Liu et al., 2023).
-The current work in hand primarily concentrates on the first application, that is using LLMs as recommender, with ChatGPT serving as the case study.
-The primary aim of this research is to enhance our understanding of “Provider Fairness within ChatGPT-based RecLLMs”.
-We focus on uncovering and understanding a range of biases – mostly underexplored in existing literature –, and examining how these biases manifest in comparison to traditional Collaborative Filtering (CF) models.
-This study looks into the dynamics of these biases, providing a detailed contrast between the two types of systems, and thereby contributing to a more nuanced understanding of the complexities inherent in modern RS.
-In particular, this research diligently identifies and analyzes inherent biases inherent in these two types of systems, such as (i) the preference for popular movies, (ii) the balance between old and recent movie recommendations, (iii) the tendency to recommend certain movie genres, (iv) the consistency/stability of recommendations over time.
-For instance, GPT-based models were found to have a penchant for suggesting recent films like those released post-2000, exemplified by their frequent recommendations of newer titles such as ”Inception” (2010) and ”The Dark Knight” (2008), compared to CF models that often favored classics like ”The Godfather” (1972) and ”Casablanca” (1942).
-Furthermore, genre dominance was evident, with GPT models showing a distinct inclination towards ‘Drama’ and ‘Comedy’, whereas CF models predominantly recommended ‘Action’ and ‘Sci-Fi’ films.
-Another crucial aspect of this research is exploring the impact of prompt engineering in a zero-shot learning setting on the performance of RecLLMs in terms of personalization, diversity, recency, and fairness.
-Zero-shot learning refers to the model’s capacity to handle tasks for which it has not been explicitly trained, relying instead on its general language understanding and adaptability.
-Here our study seeks to provide an answer to the following question: Can prompts serve as a controllability tool to steer the recommendation process in a specific direction (increase accuracy, reduce provider unfairness) based on the interests of providers or other stakeholders, and what potential harms are associated with this approach?
-Figure 1. Conceptual idea behind our work
-Example.
+Context.
+Recommender systems are integral to various large-scale internet services, benefiting consumers, producers, and other stakeholders in multi-stakeholder markets (Deldjoo et al., 2023; Abdollahpouri and Burke, 2021; Ekstrand et al., 2019; Zheng, 2019).
+The advent of generative models, particularly large language models (LLMs), holds the promise of offering better personalization experiences.
+LLMs enable conversational natural language (NL) interactions (He et al., 2023; Biancofiore et al., 2024), unlocking rich NL data like item descriptions, reviews, and queries within recommender systems (RS).
+Harnessing the general reasoning abilities of pretrained LLMs allows for addressing diverse, nuanced NL user preferences and feedback through highly personalized interactions.
+This contrasts with rigid ID-based methods that heavily rely on non-textual data.
+(Deldjoo et al., 2024a, b).
+This paper focuses on the emerging role of LLMs, specifically ChatGPT, in recommender systems and scrutinizes their biases, with a particular emphasis on “item-side fairness.”
+Item-side fairness is important to ensure diverse item groups receive fair exposure, benefiting item producers such as micro-businesses in job recommendations and promoting content related to vulnerable populations.
+Recent research shows that LLM-based recommender systems (RecLLMs), with their reliance on semantic clues and generative capabilities, can introduce unique biases not present in conventional systems (Deldjoo and Di Noia, 2024; Jiang et al., 2024; Deldjoo et al., 2023; Deldjoo, 2024; Zhang et al., 2023a).
+Therefore, it is essential to investigate how these biases manifest and impact various stakeholders.
+1.1. Background and Motivation
+At a high level, the idea of generative modeling involves creating applications that not only make decisions based on data but also generate new data by learning patterns.
+This powerful idea has enabled various applications in AI disciplines, such as image generation, text synthesis, and music composition (Min et al., 2023; Koh et al., 2024; Ferreira et al., 2023; Zhang et al., 2023b).
+Generative modeling has recently gained prominence due to advances in model paradigms such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), Variational Autoencoders (VAEs) (Kingma and Welling, 2013), Diffusion Models (Sohl-Dickstein et al., 2015; Ho et al., 2020), and Transformer-based architectures such as GPT (Wei et al., 2022a; Bubeck et al., 2023) and other LLMs.
+In the recommender system (RS) field, generative models are not entirely new and have been used in various capacities for tasks such as data synthesis and augmentation (Chae et al., 2019; Yuan et al., 2020), model regularization (Liang et al., 2018; Sachdeva et al., 2019; Ma et al., 2019), and generating complex recommendation structures (Jiang et al., 2018; Liu et al., 2021).
+These advancements fall under Deep Generative Models (DGMs), which combine traditional probabilistic models with Deep Neural Networks (DNNs).
+The core strength of DGMs lies in their ability to model and sample from the data distribution they are trained on for various inferential purposes.
+Figure 1. Conceptual idea behind experiment 1, prompt-design scenarios.
+However, these models have experienced a resurgence of interest in the RS community, similar to trends in the broader AI landscape.
+This resurgence can be partly attributed to the introduction of LLMs such as ChatGPT, which have significantly improved the natural language understanding capabilities of these systems.
+LLMs demonstrate impressive in-context learning (ICL) and few-shot generalization abilities, making them valuable assets in both direct and indirect applications within recommender systems, where our goal here is to use them as direct recommenders (shown in Figure 1).
+Notwithstanding their great success, the vast and unregulated nature of the internet data used to train LLMs raises concerns about possible biases against specific races, genders, popular brands, and other sensitive attributes that could be encoded in these networks.
+For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands.
+Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways.
+Another example could be the over-representation of content from certain geographical regions, leading to a lack of diversity in recommended media or news articles.
+Hence, the unchecked employment of these systems in commercial recommender systems may lead to unfair treatment of minority groups, reinforcing existing stereotypes, or exacerbating economic disparities.
+This research explores different biases of ChatGPT-based RecLLMs, specifically on the provider side, to understand how these biases manifest in recommendations and affect various stakeholders.
+The primary goal of this research is two-fold:
+  1. (1)
+    To enhance our understanding of the general performance of ChatGPT-based RecLLMs in terms of accuracy, provider fairness, and other nuanced evaluation dimensions (genre dominance, temporal stability, temporal freshness); and
+  2. (2)
+    To determine whether factors such as prompt design and ICL type can be leveraged in RecLLM design to enhance item fairness without compromising accuracy;
+To the best of our knowledge, the factors considered in this study are novel and have not been given enough attention in previous research.
+We believe that combining these studies into a single research effort opens up new avenues for future research in RecLLMs.
+Through two separate but extensive experiments, we introduce the following catalysts in the current work in hand:
+  • Prompt Design.
+We examine how different prompt structures and instructions affect recommendation outcomes.
+As shown in Figure 1 and Table 3, we designed seven prompt scenarios classified into three categories: accuracy-oriented, beyond-accuracy, and reasoning.
+For instance, prompts instructing the model to diversify resulted in recommendations that were temporally fresher but not significantly novel or diverse, leading to a significant reduction in accuracy (e.g., NDCG) by 50%.
+In contrast, accuracy-oriented and reasoning-focused prompts, such as Simple and Chain-of-Thought (COT), performed much better.
+  • System Roles.
+We further investigated the impact of assigning specific system roles, such as “acting as a fair recommender.” We observed that incorporating statements about fairness tends to improve system quality more effectively than including fairness directives directly within the prompt.
+  • Stability Over Time.
+We considered that external factors such as trending data or algorithmic updates might introduce randomness into the responses of GPT-based models.
+To study this, each experiment was conducted five times specifically for GPT-based models.
+We assessed the consistency of recommendations over multiple runs to ensure reliability.
+  • In-Context Learning (ICL).
+We analyzed the effect of zero-shot versus few-shot in-context learning (ICL) on provider fairness, focusing on demographic information impact on recommendations.
+We noted that incorporating user demographic data in prompts affected model biases and stereotypes.
+We structured the user profile into context, example, and demographic information parts.
+Our study highlights the nuanced effects of user profile attributes and sampling strategies on recommender systems.
+Example related to Experiment 1.
 We illustrate the impact of different “prompt design” strategies on the output of ChatGPT-based recommendation systems through a clear example, as shown in Figure 1.
@@
 To highlight the differences in recommendations based on the design of the system, we designed a total of seven prompts categorized into three distinct classes.
-These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the system’s outputs can vary significantly depending on the design strategy employed.
+These prompts were tailored to guide the recommendation system in different directions, showcasing how even with the same underlying user data, the outputs of the system can vary significantly depending on the design strategy employed.
   1. (1)
@@
 An example of recommendation under this strategy could be “The Shawshank Redemption (1994)”, accompanied by an explanation of its themes of hope and redemption that resonate with the motivational depth found in “Toy Story.”
-Promoting fairness in RecLLM: To combat popularity bias, we tested two approaches: (1) Instructional prompts: Statements like “Focus on fair recommendations, balancing popular and lesser-known movies” were integrated into various scenarios, aiming to influence RecLLM’s output.
-(2) Role-based directives: We explored directly incorporating fairness into the “system’s role”, testing its effectiveness compared to solely instructional prompts.
-This experiment seeks to identify the most effective strategies for achieving fair recommendations in RecLLM.
-In sum, our work demonstrates how strategically designing a prompt’s semantic intent and structure can impact a recommender system’s performance across accuracy, provider fairness, diversity, and temporal freshness (recency).
-Contributions. The following is a list of proposed contributions in the current work:
-  1. (1)
-    Enhanced Analysis of Prompt Design in Zero-Shot RecLMMs.
-This research provides an in-depth study into how the design of prompts (in terms of intent, structure, and system role) affects the performance of ChatGPT-based recommendation systems.
-It investigates how different prompt design impacts factors such as accuracy, fairness, diversity, and temporal freshness in recommendations.
-This study involves experimental designs that compare various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and analyze their impacts on the recommendation of niche versus popular content, as well as the balance between old and recent movies.
-  2. (2)
-    Identification and Analysis of Different Biases in Zero-Shot RecLLMs.
-This work identifies and analyzes numerous biases that may be present in RecLLMs such as ChatGPT, thus laying the groundwork for investigating bias mitigation strategies.
-Key biases identified under scrutiny include popularity bias, genre preference bias, and temporal (recency) bias.
-We study how these biases uniquely manifest in LLM-based systems versus traditional Collaborative Filtering (CF) models.
-  3. (3)
-    Stability Analysis in Zero-Shot RecLLMs: This world also studies the stability of recommendations provided by LLM-based systems over time.
-Stability in this context refers to the consistency of recommendations given similar input conditions at different times.
-Our study analyzes and discusses factors that might impact stability (besides the parameters controlled by LLMs), such as changes in underlying data, model updates, and prompt variations.
-It would also explore the implications of stability for user satisfaction and trust in the system in real-life deployments.
+1.2. Contributions.
+We designed two carefully-structured experiments focusing on classical top-k recommendation and sequential in-context learning.
+These experiments involved thousands of API calls to gather the findings presented in this work.
+The first experiment investigates how different prompt design strategies impact recommendation accuracy, provider fairness, as well as genre dominance and temporal freshness in classical top-k recommendations.
+The second experiment evaluates the effectiveness and fairness of RecLLMs in sequential in-context learning tasks, comparing zero-shot and few-shot learning scenarios across different datasets.
+This work makes several contributions to the field of recommender systems, as listed below:
+  • Enhanced Analysis of Prompt Design in Zero-Shot RecLLMs: We provide an in-depth study of how different prompt designs (in terms of intent, structure, and system role) affect the performance of ChatGPT-based recommendation systems.
+This includes an analysis of various prompt strategies (e.g., accuracy-oriented, beyond-accuracy-oriented, explanation-oriented) and their impacts on recommendation personalization, diversity, recency, and fairness.
+  • Identification and Analysis of Biases in Zero-Shot RecLLMs: Our research identifies and analyzes numerous biases present in RecLLMs, such as item fairness, genre preference bias, and temporal (recency) bias.
+We compare these biases with traditional CF models, providing insights into how they uniquely manifest in LLM-based systems;
+  • Stability Analysis in Zero-Shot RecLLMs: We study the stability and consistency of recommendations provided by LLM-based systems over time.
+Specifically, we examine variations caused by changes in underlying data, model updates, and their implications on the quality of recommendations obtained by RecLLMs;
+  • Generalization of Fairness Principles in Sequential Recommendations: We evaluate the effectiveness of RecLLMs in sequential recommendation tasks across different datasets, focusing on zero-shot versus few-shot learning.
+This analysis provides insights into the adaptability and robustness of these models in maintaining fairness across varied recommendation contexts.
+  • Impact of User Profile Attributes and Demographic Information on Fairness and Accuracy: We explore the influence of user profile attributes (historical interaction sampling) and the inclusion of demographic information (e.g., gender, age-group) on recommendation fairness and accuracy, providing new insights into how revealing such information can affect model biases and stereotypes.
+These contributions collectively enhance our understanding of prompt-based recommendation systems using LLMs, offering new perspectives on their biases, fairness, and stability.
+Our findings highlight the potential and challenges of integrating LLMs into recommender systems, paving the way for future research and improvements in this domain.
 The structure of the paper is organized as follows.
@@
 Section 3 is dedicated to presenting a thorough evaluation of ChatGPT.
-This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1), the repeated experiments for stability analysis (Section 3.2), an understanding of the system’s role in ChatGPT-based RecLLM (Section 3.3), and the promotion of fairness (Section 3.4).
-Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.5).
+This is this research’s core contribution which we present through the suite of goal-oriented prompts’ (Section 3.1.1), the repeated experiments for stability analysis (Section 3.1.2), an understanding of the system role in ChatGPT-based RecLLM (Section 3.1.3), and the promotion of fairness (Section 3.1.4).
+Additionally, we explore the explicit versus implicit scenario, which is indicated in prompt structuring (Section 3.1.5).
 Following this, Section 4 outlines the experimental setup employed in our study.
@@
 2.1. Fairness in Recommender Systems
-Fairness has emerged as a pivotal topic within the AI community in the last few years, arguably becoming the most scrutinized aspect across various branches of trustworthy AI, including fairness, security, privacy, and explainability.
-In RS, fairness has garnered significant attention due to the multi-stakeholder nature of these systems (Deldjoo et al., 2023; Ekstrand et al., 2019).
+Fairness has become a pivotal topic in AI, gaining scrutiny across branches of trustworthy AI, including security, privacy, and explainability.
+Fairness in recommender systems has gained significant attention due to their multi-stakeholder nature (Deldjoo et al., 2023; Ekstrand et al., 2019).
 Unfairness, even in its minimal form, can adversely impact various stakeholders, including consumers, producers, system designers, supply chains, and even the environment – the latter are often referred to as ‘side-stakeholder’.
-The body of literature on fairness in RS is diverse, encompassing multiple perspectives and dimensions.
-Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), have introduced various taxonomies to categorize these diverse aspects into several orthogonal or partially-orthogonal dimensions.
+The body of literature on fairness in RS is diverse, covering multiple perspectives and dimensions.
+Recent surveys (Ekstrand et al., 2019; Deldjoo et al., 2023; Amigó et al., 2023), categorize these aspects into several dimensions, either orthogonal or partially orthogonal.
 Key dimensions recognized in the literature include the main stakeholder in question (e.g., consumer vs.
@@
 These dimensions offer a comprehensive view of the fairness landscape in RS, as detailed in these surveys.
-In the context of the present study, and with regard to positioning this work within the literature of fair Recommender Systems (fair-RS), we have introduced Table 1.
+To position this study within the literature of fair recommender systems, we introduced Table 1.
 This table is designed to categorize existing literature along two principal dimensions: the stakeholder in question (consumer vs.
@@
 This distinction is crucial because, while RecLLM, such as those based on GPT-like architectures, promise to significantly advance RS landscape with more nuanced and personalized recommendations, they also raise concerns about inherent biases.
-The vast and unregulated nature of internet data used for training LLMs raises concerns about inherent biases against specific races, genders, popular brands, and other sensitive attributes.
-For example, if an LLM is predominantly trained on data from popular e-commerce sites, it might disproportionately recommend products from more recognized brands, overlooking niche or emerging brands.
-Similarly, biases in language around gender or race could skew recommendations in subtle but impactful ways.
-Therefore, accurately measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies, ensuring that advancements in RecLLM do not amplify existing inequalities.
-This is the key goal of the current work in hand.
+The extensive and unregulated nature of internet data used to train LLMs raises concerns about biases against specific races, genders, and other sensitive attributes.
+For example, if an LLM is trained on e-commerce data where men’s products are described more positively than women’s products, it might skew recommendations towards male-oriented items.
+Therefore, measuring these biases is a crucial initial step, and forms the key goal of our current work, in developing effective mitigation strategies.
+We can briefly review the research on fairness in recommendation systems according to Table 1 as follows:
   • Traditional RS.
-This column lists studies that have focused on traditional methods of recommender systems.
+This column lists studies that have focused on traditional methods of recommender systems (Deldjoo et al., 2018).
 They primarily rely on CF algorithms (possibly using side information of users and items) without the advanced natural language processing capabilities of LLMs.
@@
   • RecLLM: This column spotlights the burgeoning research domain that merges language models (LMs) with RS representing a significant shift towards harnessing advanced NLP techniques to enhance the accuracy and relevance of recommendations.
-These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023a; Zhang et al., 2023).
-Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023a; Zhang et al., 2023), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
+These studies explore the use of various LMs, including BERT-based models (Shen et al., 2023) and recent LLMs such as GPT-like architectures (Li et al., 2023; Zhang et al., 2023a).
+Additionally, beyond the scale of LMs, the target tasks within RS–such as classical recommendation (top-$k$ ranking (Li et al., 2023; Zhang et al., 2023a), sequential), conversational RS (Shen et al., 2023), explanation generation, multi-modal recommendations–provide dimensions that could be used for further categorizing these works.
 Table 1. Research landscape in Recommender Systems focusing on Consumer and Producer fairness with respect to Age, Gender, Other, and Others
@@
 Consumer Fairness Activity (Li et al., 2021; Hao et al., 2021; Xiao et al., 2020; Rahmani et al., 2022b)
-Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023)
+Demographics (Deldjoo et al., 2021a, b; Wu et al., 2021b; Weydemann et al., 2019; Farnadi et al., 2018) (Shen et al., 2023; Zhang et al., 2023a; Deldjoo and Di Noia, 2024; Deldjoo, 2024)
 Merits (Sühr et al., 2021; Gómez et al., 2021)
 Others (Wan et al., 2020; Lin et al., 2021)
-Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023a)
+Producer Fairness Popularity (Dong et al., 2021; da Silva et al., 2021; Ge et al., 2021; Zhu et al., 2021) (Li et al., 2023)
 Demographics (Kırnap et al., 2021; Boratto et al., 2021)
@@
 Stakeholder.
-As mentioned earlier, a dominant aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides.
+As mentioned earlier, a major aspect that can be utilized to classify almost all literature on group fairness is the market side focus—whether they concentrate on a single-side market (defined either by consumers or providers) or on both sides (Deldjoo et al., 2022; Rahmani et al., 2024; Naghiaei et al., 2022).
 Within each of these segments, as you can observe, we can further categorize the literature based on which sensitive attribute groups are defined.
@@
 It includes studies that focus on ensuring fairness among consumers of the recommender system based on these attributes.
-For instance, ensuring that recommendations are not biased towards a particular demographic group.
+For instance, ensuring that recommendations are not biased towards a particular demographic consumer group.
   • Producer Fairness.
@@
 The study by Zhang et al.
-(2023) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness.
+(2023a) evaluates the consumer fairness side of zero-shot GPT recommendations, focusing on a variety of consumer demographic attributes but not addressing producer-side fairness.
 Their work introduces a novel benchmark, FaiRLLM, for evaluating the fairness of RecLLM, highlighting ChatGPT’s biases towards certain sensitive user attributes in music and movie recommendations.
+The work by Deldjoo et al.
+(Deldjoo and Di Noia, 2024; Deldjoo, 2024) address the shortcomings by scrutinizing whether changes in recommendations, due to the inclusion of sensitive attributes, result in unfairness.
+It also explores better ways to normalize and improve the fairness of these models.
 Conversely, the work by Li et al.
-(2023a) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation.
-This study investigates ChatGPT’s performance in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection.
-Similarly, research such as that by Di Palma et al.
-(2023) examines the personalization and popularity bias in RecLLMs, particularly in the context of ChatGPT and other LLMs.
+(2023) aligns more closely with ours, focusing on producer unfairness however within the specific domain of news recommendation.
+This study investigates performance of ChatGPT in news recommendation, exploring aspects like personalization, provider fairness, and fake news detection.
 However, these studies do not address the intricacies of prompt engineering in RecLLMs, nor do they comprehensively address the various forms of biases studies in the current study.
-Our research goes beyond examining provider fairness based on popularity bias and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.
+Our research goes beyond examining provider fairness based on item unfairness and also considers other potential harms, such as the recency of recommended items and the stability of recommendations—aspects not explored in the aforementioned studies.
 2.2. Leveraging Pre-trained LMs and Prompting for Recommender Systems
@@
 This method, including knowledge distillation techniques, shows potential in intra-task prompt distillation and cross-task prompt transfer, contributing to the efficiency and effectiveness of LLM-based recommendation models.
+In recent surveys (Deldjoo et al., 2024a, b), the wide area of generative models has been extensively explained.
+These surveys highlight the potential of generative models in transforming traditional RS by leveraging their capabilities in understanding and generating complex data distributions.
+For instance, they have shown how generative models can effectively handle multimodal data (text, images, and videos), thereby enhancing the richness and accuracy of recommendations.
+These surveys also shed light on the emergent reasoning abilities of LLMs, such as in-context learning, which allows models to adapt to new tasks with minimal additional training.
+Moreover, they discuss the application of retrieval-augmented generation (RAG) techniques, which combine information retrieval with generative modeling to produce contextually relevant recommendations.
+Overall, these insights underscore the significant advancements and ongoing challenges in utilizing generative models for RS.
 3. Evaluation of ChatGPT-based RecLLM
-Our work focuses on evaluating ChatGPT as a zero-shot recommender in the movie domain, aiming to extend beyond previous research by delving more deeply into biases and potential harms caused by such systems.
-Given the extensive nature of our experiments, we chose to concentrate on the single domain of movie recommendations.
-In this context, our evaluation methodology for ChatGPT was comprehensive and multifaceted.
-We conducted extensive experiments across a multi-dimensional testbed to assess the robustness and adaptability of the system.
-This evaluation particularly emphasized understanding the role and impact of ChatGPT in three key areas:
-  1. (1)
-    Goal-oriented prompt,
-  2. (2)
-    System role,
-  3. (3)
-    Promoting fairness, as well as
-  4. (4)
-    Implcit vs. explicit scenario
-These dimensions were varied to observe their individual and combined effects on the quality of the recommendations provided.
-Please note that since the primary objective of this study is centered on (provider) fairness, the emphasis on fairness has been integrated either as a distinct scenario within the goal-oriented prompts or within the system role itself, and the impact is evaluated.
-3.1. Goal-oriented prompts
+We designed two experiments to evaluate the performance and bias of ChatGPT-based RecLLMs.
+  • Experiment 1. Examining prompt design strategies in classical top-$k$ recommendation (cf. Section 3.1), and
+  • Experiment 2. Studying generalization of findings across datasets, and learning strategies (zero vs. few-shot ICL) in sequential recommendation task (cf. Section 3.2)
+Table 2 provides a summary of these experiments.
+Table 2. Summary of Experiments
+Exps. Aim Approach Evaluation Metrics
+Exp 1 (i) Investigate prompt design strategies in classical top-K recommendations (ii) Analyze stability of results over different RecLLM runs, (iii) Look at new content-related metrics (genre diversity, temporal freshness) Explore prompt formulations (accuracy, diversity, fairness), system role directive, specific information in prompt in the movie domain Accuracy (top-K precision and recall), beyond-accuracy metrics (diversity, novelty), fairness metrics, genre diversity, temporal freshness
+Exp 2 (i) Assess generalization of fairness across datasets and beyond zero-shot learning to ICL, (ii) Look at the impact of user profile construction and demographic information revealing Use movie and music dataset, evaluate few-shot and zero-shot ICL in sequential recommendation Accuracy, provider fairness, catalogue coverage
+3.1. Experiment 1: Examining Prompt Design Strategies in Classical Top-K Recommendations
+In the first experiment, we investigate prompt design strategies in a classical top-$K$ recommendation setting.
+We use an LLM (ChatGPT) as a zero-shot recommender, querying it with various prompt formulations to obtain recommendations.
+Our goal is to understand how much the recommendation outcome (performance) can be influenced by prompts carrying different objectives (e.g., relevance, increasing diversity, or motivating reasoning).
+ur goal is to determine if a one-to-one mapping exists between the desired natural language query (e.g., enhance diversity) and the actual diversity of the model recommendations.
+To achieve the goals of this paper, we examine the relevance and item fairness (including diversity) of the recommendations.
+Additionally, we explore new aspects of recommendation, such as content diversity and freshness, including genre diversity and temporal freshness.
+Table 3. Overview of prompt scenarios designed for the experiment 1.
+Scenario Description and Prompt
+Personalization Focused
+S1 - Simple Basic recommendations without additional context.
+Prompt: ”Recommend 10 movies that the user will likely enjoy.”
+S2 - Genre-focused Recommendations focusing on genres and themes similar to the user’s past favorites.
+Prompt: ”Recommend 10 movies that the user will likely enjoy, particularly focusing on genres and themes similar to their past favorites.”
+S3 - Rating-focused Incorporation of the user’s explicit ratings into the recommendations.
+Prompt: ”Recommend 10 movies the user will likely enjoy, taking into account both their favorite genres and past movie ratings.”
+Beyond-Accuracy Focused
+S4 - Diversify Recommendations Suggesting lesser-known films to diversify the user’s experience.
+Prompt: ”Suggest 10 high-quality, lesser-known films that diverge from mainstream blockbusters, yet align with the user’s tastes.”
+S5 - Surprise Offering unexpected recommendations for exploring new preferences.
+Prompt: ”Offer 10 unique and unexpected movie recommendations aimed at broadening the user’s cinematic horizons beyond their usual preferences.”
+Reasoning-Explanation Focused
+S6 - Motivate Reasoning Providing reasoning for each recommendation to enhance transparency.
+Prompt: ”Provide 10 carefully selected movie recommendations, each accompanied by a rationale explaining its suitability for the user’s preferences.”
+S7 - Chain of Thought (COT) Engaging in a logical reasoning process to arrive at the recommendations.
+Prompt: ”Let’s think this through: What would be 10 great movie recommendations for this user and why?”
+3.1.1. Goal-oriented prompts
 Seven distinct prompt scenarios were designed, categorized into three classes focused on personalization, beyond-accuracy metrics, and reasoning.
@@
 The scenarios are detailed in Table 3 and further exemplified in Section 1.
-Table 2. Overview of Prompt Scenarios.
-Scenario Description and Prompt
-Personalization Focused
-S1 - Simple
-Basic recommendations without additional context.
-Prompt: ”Recommend 10 movies that the user will likely enjoy.”
-S2 - Genre-focused
-Recommendations focusing on genres and themes similar to the user’s past favorites.
-Prompt: ”Recommend 10 movies that the user will likely enjoy, particularly focusing on genres and themes similar to their past favorites.”
-S3 - Rating-focused
-Incorporation of the user’s explicit ratings into the recommendations.
-Prompt: ”Recommend 10 movies the user will likely enjoy, taking into account both their favorite genres and past movie ratings.”
-Beyond-Accuracy Focused
-S4 - Diversify Recommendations
-Suggesting lesser-known films to diversify the user’s experience.
-Prompt: ”Suggest 10 high-quality, lesser-known films that diverge from mainstream blockbusters, yet align with the user’s tastes.”
-S5 - Surprise
-Offering unexpected recommendations for exploring new preferences.
-Prompt: ”Offer 10 unique and unexpected movie recommendations aimed at broadening the user’s cinematic horizons beyond their usual preferences.”
-Reasoning-Explanation Focused
-S6 - Motivate Reasoning
-Providing reasoning for each recommendation to enhance transparency.
-Prompt: ”Provide 10 carefully selected movie recommendations, each accompanied by a rationale explaining its suitability for the user’s preferences.”
-S7 - Chain of Thought (COT)
-Engaging in a logical reasoning process to arrive at the recommendations.
-Prompt: ”Let’s think this through: What would be 10 great movie recommendations for this user and why?”
 In particular, scenarios S1 to S3 are focused on core personalization aspects, encompassing basic recommendations without additional context (S1), genre preferences (S2), and the incorporation of explicit user ratings (S3).
@@
 Scenarios S6 and S7 are explanation-motivation oriented, where S6 seeks to provide reasoning for each recommendation to enhance transparency, and S7 involves a logical ste-by-step reasoning process to arrive at the recommendations.
-These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022; Wang et al., 2022; Chu et al., 2023).
-Note.
+These scenarios compose a competitive and insightful set of prompts, some of which such as COT, have been successfully tested in other ML disciplines (Wei et al., 2022b; Wang et al., 2022b; Chu et al., 2023).
+[^1] Note.
 The CF baselines used in this work operate in an implicit setting.
@@
 Scenario ‘Rating-focused’ S3 is the only exception, included solely for completeness.
-This is discussed in Section 3.5.
-3.2. Repeated Experiment for the Stability of the Analysis
+This is discussed in Section 3.1.5.
+3.1.2. Repeated Experiment for the Stability of the Analysis
 The experiments were conducted at a temperature setting of 0.0, aiming for predictable responses.
-However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in GPT’s behavior.
+However, we observed variations in the recommendations, possibly due to factors like trending data, recent updates in the database, or changes in the behavior of ChatGPT.
 This suggests that the recommendation system may interpret the same input differently in each iteration, independent of the temperature setting.
 To assess the robustness and reliability of our system, we repeatedly conducted recommendation queries for each user, five times each, to examine if outputs vary despite consistent input.
-We further adopted a “bootstrapping sampling strategy” during the evaluation stage, as detailed in Section 4.3.
+We further adopted a bootstrapping sampling strategy during the evaluation stage, as detailed in Section 4.1.3.
 This method is particularly effective in revealing the reliability of the results, even when dealing with datasets that are small or may not perfectly represent the entire population.
-This thorough evaluation approach, therefore, enhances our confidence in the system’s performance across various potential real-world scenarios.
-3.3. Understanding ChatGPT’s “System” Role
+This thorough evaluation approach, therefore, enhances our confidence in the performance of the system across various potential real-world scenarios.
+3.1.3. Understanding the impact of “System” Role in ChatGPT
 In this study, we sought to understand the impact of different system roles assigned to ChatGPT on recommendation outcomes.
 We focused on assessing whether responses significantly differ when specific “system” instructions are attached, and how these variations influence the functionality of a recommender system.
-Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system’s operations versus incorporating it in the prompt structure.
+Specifically, we aimed to determine the effectiveness of embedding fairness directly into the system role versus incorporating it in the prompt itself.
 Table 4 summarizes the roles assigned to ChatGPT during our experiments:
@@
 Role ID Description
-R0 - No Role
-Direct user-driven prompts without system context.
-R1 - System Role as Recommender
-System-centric role focusing purely on user information.
-R2 - System Role as Fair Recommender
-System role with an explicit fairness objective.
-The roles explored provide insight into how ChatGPT’s framing, in terms of its responsibilities and objectives, shapes its behavior in making recommendations.
-This analysis is crucial for comprehending the subtleties of human-AI interaction within diverse contexts and for distinct purposes, especially in how the system interprets and adheres to varying directives.
-3.4. Promoting Fairness
+R0 - No Role Direct user-driven prompts without system context.
+R1 - System Role as Recommender System-centric role focusing purely on user information.
+R2 - System Role as Fair Recommender System role with an explicit fairness objective.
+3.1.4. Fairness Emphasis
 We also aim to understand whether integrating a fairness statement in the system or in the prompt itself is more effective.
@@
 The latter prompt is combining Scenario (S1/E1).
-3.5. Explicit vs. Implicit Scenario
-This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the recommender system’s quality, based on the premise that including such information could improve recommendations.
+3.1.5. Explicit vs. Implicit Scenario
+This test evaluated whether displaying the movies a user has rated, as indicated in {user_movies_string}, enhances the quality of RS, based on the premise that including such information could improve recommendations.
 Scenarios S1 to S6, except S3, were conducted implicitly, without user ratings in the prompts, mirroring the implicit mode of the CF baseline experiments.
@@
 The Matrix (Genres: Action/Sci-Fi, Rating: 5/5), Inception (Genres: Action/Thriller, Rating: 4.5/5)
-The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendation more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
+The primary hypothesis under investigation is whether the inclusion of explicit ratings enables the system to tailor recommendations more closely to the user’s demonstrated preferences, thereby enhancing the personalization of the recommendation process.
+Figure 2. Sequential in-context learning for various scenarios explored in Experiment 2 of this research.
+3.2. Experiment 2: Sequential In-Context Learning
+The task of recommending suitable movies/music to users can be modeled as a sequential ranking problem where the historical interactions of a user are utilized to predict future preferences.
+In this context, the goal is to rank candidate items $C=\{i_{j}\}_{j=1}^{m}$ such that the most relevant items for the user appear at the top.
+Given a user’s interaction history $H=\{i_{1},i_{2},\ldots,i_{n}\}$, to leverage it for generating recommendations, we opt to divide it into two main parts, as displayed in Figure 2:
+  1. (1)
+    Context part.
+This part is sampled from the entire interaction history of the user to provide general context about the user’s interests, passion, and preference (toward music and movies).
+We use two strategies for sampling:
+    • Frequent/Highest-rated: Selects the most enjoyed items by the user, based on the ratings provided (for movies) or frequency of consumption (for music);
+    • Recency-Focused: Utilizes a weighted average method to emphasize recently consumed items.
+This strategy selects items that have not only been liked by the user but also have been recently listened to or watched, ensuring relevance and timeliness in the recommendations.
+  2. (2)
+    Example part.
+This part is taken from recent interactions to guide specific next recommendations.
+It directly determines the next items to be consumed in sequential tasks and is used in ICL-$1$, and ICL-$2$, which denote the few-shot learning scenarios, using one or two examples.
+Based on this division, we explore different prompting strategies, as outlined in the following.
+We note that Experiment 2 has been inspired by Hou et al.
+(2024) and extends it by exploring the above profile investigation, and evaluation going beyond pure accuracy.
+Note.
+In the prompting scenarios outlined below, we optionally incorporate “user demographic information” (such as gender, or age category) into the prompts to examine how revealing this information might influence model biases and stereotypes.
+Specifically, where we have {is demographic} information, it is used in statements like ‘‘The user is female’’, ‘‘The user is young’’.
+Details of the results of comparisons with and without demographic information on fairness are provided in Table 10.
+Sequential Zero-Shot Prompting
+This method only uses the context part of the user interaction history
+Example:
+The user {is female, and} has watched the following movies:
+- "From Russia with Love (1963)" in genre(s) Action with rating 5
+- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 4
+- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
+- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
+- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
+This selection reflects the user’s movie preferences.
+What would be the top-1 suitable next recommendation?
+as stated earlier, {is female} is essentially denoted with {is demographic} below, which is tested in two scenarios: included and omitted.
+Sequential In-Context Learning (ICL-1)
+This approach takes the contest part of the preivous part, but taked the xample part of the interaction and from that take the last interatcion to guide the model how the recommendation.
+The examples are derived from recent interactions.
+Example:
+The user {is demographic, and} has watched the following movies:
+- "From Russia with Love (1963)" in genre(s) Action with rating 5
+- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 3
+- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
+- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
+- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
+This selection reflects the user’s movie preferences.
+Given the user has recently watched the following movies in order:
+1. "Face/Off (1997)" in genre(s) Action|Sci-Fi|Thriller with rating 4
+2. "Bringing Out the Dead (1999)" in genre(s) Drama|Horror with rating 5
+3. "Sixth Sense, The (1999)" in genre(s) Thriller with rating 5
+4. "Austin Powers: The Spy Who Shagged Me (1999)" in genre(s) Comedy with rating 4
+5. "Arlington Road (1999)" in genre(s) Thriller with rating 4
+You should recommend:
+Recommendation 1: "Bowfinger (1999)" in genre(s) Comedy with rating 3
+What would be the top-1 suitable next recommendation after the above movies?
+Sequential In-Context Learning (ICL-2)
+Similar to ICL-1 but with a bigger selection of examples equal to 2. The examples are chosen to provide more specific guidance on generating recommendations based on recent user interactions.
+Example:
+The user {is demographic, and} has watched the following movies:
+- "From Russia with Love (1963)" in genre(s) Action with rating 5
+- "Star Trek IV: The Voyage Home (1986)" in genre(s) Action|Adventure|Sci-Fi with rating 3
+- "Planet of the Apes (1968)" in genre(s) Action|Sci-Fi with rating 3
+- "Star Wars: Episode I - The Phantom Menace (1999)" in genre(s) Action|Adventure|Fantasy|Sci-Fi with rating 4
+- "Final Conflict, The (a.k.a. Omen III: The Final Conflict) (1981)" in genre(s) Horror with rating 3
+This selection reflects the user’s movie preferences.
+Given the user has recently watched the following movies in order:
+1. "Strange Days (1995)" in genre(s) Action|Crime|Sci-Fi with rating 3
+2. "Face/Off (1997)" in genre(s) Action|Sci-Fi|Thriller with rating 4
+3. "Bringing Out the Dead (1999)" in genre(s) Drama|Horror with rating 5
+4. "Sixth Sense, The (1999)" in genre(s) Thriller with rating 5
+5. "Austin Powers: The Spy Who Shagged Me (1999)" in genre(s) Comedy with rating 4
+You should recommend:
+Recommendation 1: "Bowfinger (1999)" in genre(s) Comedy with rating 3
+Recommendation 2: "Arlington Road (1999)" in genre(s) Thriller with rating 4
+What would be the top-1 suitable next recommendation after the above movies?
+Recency-Focused Zero-shot and ICL-1 and ICL-2
+We replicate all previous prompting scenarios but modify the “context part” to include the most recent interactions.
+This approach indicates to the ChatGPT recommender system that the best recommendations are based on the user’s recent consumption.
+The next sections outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
 4. Experimental Setup
-This section outlines experiments conducted to assess various reproducibility aspects, including the obtained results and our observations.
-4.1. Datasets and splitting
-In this study, we utilized the MovieLens-Latest-Small dataset provided by the GroupLens group, [^1] https://grouplens.org/datasets/movielens/ which contains the most recent versions of movies.
-It is important to note that we intentionally focused our research within the domain of movies.
-This decision was made due to the potential varying capabilities of GPT-based models across different domains.
-Additionally, conducting extensive experiments with GPT models is time and resource-intensive, which is why we opted to work with this specific dataset.
-We acknowledge that there is potential for future exploration in other domains, and we leave this avenue open for future research.
-Detailed dataset parameters, including $\frac{R}{U}$ (ratings per user) and $\frac{R}{I}$ (ratings per item), resulting in a higher density $\frac{R}{U\times I}$, are presented in Table 4 for reference.
-Our setting involves a train-validation-test split for the data with ratios 80%, 10%, and 10%, respectively.
-4.2. Evaluation Metrics
-Our study employs a diverse range of metrics to assess the performance of our recommendation system.
-These metrics are categorized into two primary classes: Accuracy, Beyond-accuracy and Provider Fairness.
-Accuracy Metrics. These metrics measure the quality of the top-$k$ ranking in the recommendation list.
-  • NDCG@k. A measure of ranking quality, particularly the placement of relevant items.
+In this section, we detail the experimental setup for two experiments, including the datasets, evaluation metrics, bootstrapping methods, baselines employed, and hyperparameter tuning.
+4.1. Experiment 1. Classical Top-$k$ Recommendation
+4.1.1. Datasets and Setup
+Experiment 1 used the MovieLens-Latest-Small dataset from the GroupLens group.
+[^2] https://grouplens.org/datasets/movielens/ This dataset comprises the latest movie entries and was used in its entirety to carry out experiments in part 1 of our study.
+The experimental design required conducting API calls for each user, resulting in a total of 610 API calls (equal to the number of users) per individual recommendation setting.
+Given the scope—610 users, 7 different prompts or prompt scenarios (cf.
+Section 3.1.1), 2 fairness statements in the prompt (cf.
+Section 3.1.4), and 2 system roles (cf.
+Section 3.1.3)—we generated thousands of scenarios or API calls.
+Due to the computational demands, we limited our experiments in part 1 to a single dataset, focused on the classical top-$k$ recommendation task.
+The detailed statistics of the dataset are provided in Table 5, where we implemented a data split of 80% training, 10% validation, and 10% testing.
+The validation dataset is used solely for CF baseline hyperparameter tuning, while both RecLLMs and the CF baseline utilize the same training and testing datasets to conduct recommendations.
+4.1.2. Evaluation Metrics
+The first experiment employs a diverse range of metrics to assess the performance of different RS.
+These metrics are categorized into two primary classes: accuracy metrics, and provider fairness/diversity.
+Additionally, we incorporate new content-centric metrics such as temporal freshness and genre dominance/tendancy to explore insights into potential biases and tendencies of RecLMMs compared to mainstream CF models.
+Accuracy Metrics.
+These conventional metrics measure the quality of the top-$k$ recommendation or the personalization of recommendations
+  • NDCG@k: Measures the ranking quality by evaluating the placement of relevant items within the top-$k$ positions.
   • Recall@k. The proportion of relevant items that are successfully recommended.
 High scores in these metrics indicate more accurate and relevant recommendations.
-Beyond-Accuracy and Provider Fairness Metrics.
-Beyond conventional accuracy metrics, these metrics evaluate the diversity, and fairness of the recommendations w.r.t provider perspective.
+Provider Fairness/Diversity.
+Beyond conventional accuracy metrics, these metrics evaluate the fairness of the recommendations from the provider perspective.
 They assess how well the system introduces new and varied content and how equitably it treats different item providers.
-  • Gini Index: Measures the inequality in the distribution of item recommendations.
+  • Gini Index (Deldjoo et al., 2021b): Measures the inequality in the distribution of item recommendations.
 A lower Gini Index indicates a more equitable distribution among items.
-The Gini Index is beneficial for identifying monopolization of recommendations by a few items, but it may not always capture the full picture of item diversity.
-  • HHI (Herfindahl-Hirschman Index): Assesses the concentration of recommendations among items.
-A lower HHI value suggests a healthier, more diverse recommendation landscape.
-This index is particularly useful for detecting dominance by a small number of items, though it might overlook the distribution nuances among less recommended items.
-  • Entropy: Reflects the diversity and unpredictability of recommendations.
-Higher entropy values indicate greater diversity.
-Entropy is advantageous for understanding the overall unpredictability of recommendations but may not directly indicate the fairness of item exposure.
-Each of these metrics has its own advantages and limitations, making them collectively essential for a well-rounded evaluation.
-They help in understanding not just the effectiveness of the recommendation system in terms of user satisfaction (accuracy) but also its robustness in promoting diverse and fair content distribution.
-4.3. Bootstrapping Sampling Strategy
-In our evaluation stage, we employed a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement.
+The Gini index is calculated as follows:
+    $\text{Gini}=\frac{\sum_{i=1}^{n}(2i-n-1)x_{i}}{n\sum_{i=1}^{n}x_{i}}$
+    where $x_{i}$ represents the number of times item $i$ is recommended, sorted in ascending order, and $n$ is the total number of items.
+The Gini index becomes 0 when all items are recommended an equal number of times, indicating perfect equality.
+Mathematically, if all $x_{i}$ are equal, then:
+    $x_{i}=\frac{\sum_{i=1}^{n}x_{i}}{n}\quad\forall i\implies\text{Gini}=0$
+    The Gini index becomes 1 when one item is recommended exclusively, indicating maximum inequality.
+In our work, we interpret the Gini index to measure item fairness, where lower Gini values indicate higher (better) item fairness.
+  • HHI (Herfindahl-Hirschman Index) (of Justice, 2018; Wang et al., 2022a): This metric Assesses the concentration of recommendations among items. It is calculated as follows:
+    $\text{HHI}=\sum_{i=1}^{n}\left(\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}\right)^{2}$
+    The HHI value ranges from $1/n$ to 1, where $n$ is the total number of items.
+The HHI is $1/n$ when all items are recommended equally, indicating maximum fairness, and it is 1 when one item is recommended exclusively, indicating no fairness.
+Here, lower values of the HHI are considered indicative of better item fairness.
+  • Entropy: Reflects the diversity and unpredictability of recommendations. Higher entropy values indicate greater diversity. The Entropy is calculated as follows:
+    $\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i})$
+    where $p_{i}=\frac{x_{i}}{\sum_{j=1}^{n}x_{j}}$ and $p_{i}>0$.
+The entropy value ranges from 0 to $\log(n)$, where 0 indicates no diversity (one item recommended exclusively) and $\log(n)$ indicates maximum diversity (all items recommended equally).
+Higher values of Entropy are considered indicative of better item fairness.
+While the above metrics are used to measure item fairness based on item exposure, they also serve as indicators of how recommendations produce diverse outcomes.
+4.1.3. Bootstrapping Sampling Strategy
+In experiment 1, for the evaluation, we additionally employ a bootstrapping sampling strategy, a statistical technique where 1000 samples were generated by repeatedly resampling our dataset with replacement (Deldjoo, 2023).
 This method was chosen to enhance the robustness and reliability of our analysis, particularly in measuring the performance metrics of our recommendation system, such as NDCG, Recall, and Precision.
 Through this approach, we computed means for each metric across all bootstrap samples, then calculated their averages and 95% confidence intervals.
-This strategy is implemented for understanding the variability and confidence in our metrics, ensuring our evaluation reflects the system’s likely real-world performance.
-It is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
-4.4. CF Baselines.
+This strategy is implemented for understanding the variability and confidence in our metrics, to ensure our evaluation reflects the likeness of the system performance on real-world performance.
+This is especially beneficial in situations with small or non-representative samples, allowing for more reliable population inferences.
+4.1.4. CF Baselines.
 The choice of models and the corresponding hyperparameter optimization process is outlined as follows.
@@
 MovieLens Latest Small 610 9,724 100,836 165.30 10.36 98.3% 0.715 0.603
-4.4.1. Models.
+Models.
 We use a suite of competitive, CF recommendation models as baseline ranking models in our post-processing approach, as summarized below.
@@
 Additionally, the TopPop method is included as a non-informative baseline, as an essential baseline to provide a case where the outcomes are heavily biased towards popular items.
-4.4.2. Hyperparameter Tuning.
+Hyperparameter Tuning.
 In our research, the primary focus is on the hyperparameter tuning of various collaborative filtering (CF) recommendation models, selected as baselines.
@@
 Hyperparameter tuning is an essential step in optimizing the performance of these models.
-We employ a greedy search strategy to identify the best configurations based on their performance on the validation set.
+We used a bootstrapping sampling strategy, generating 1000 samples by repeatedly resampling our dataset with replacement.
 The following is a summary of the hyperparameter tuning conducted for each model:
@@
 The total number of hyper-parameters is 14, encompassing a total of 223 experimental cases.
-This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our resear
-5. Results
-In this section, we present insights gained from the experimental validation of our system.
-We have divided our analysis into four key areas, each focusing on different aspects of performance: (i) personalization (cf.
-Section 5.1), (iii) stability of recommendations over different runs (cf.
-Section 5.2), (ii) provider fairness and diversity of recommendations (cf.
-Section 5.3), and finally (iv) the temporal freshness (or recency) of recommendations (cf.
-Section 5.4).
-We also provide an interesting discussion on the genre dominance/bias of recommendation models employed in Section 5.5.
-We approach this section by formulating a number of the hypotheses for each part, afterward showcasing the results, and providing the associated analysis.
-Finally, we will provide answers to these hypotheses.
-5.1. Personalization of Recommendations
+This approach ensures that we effectively explore a wide range of hyperparameter configurations, particularly for models operating in implicit feedback scenarios, which are crucial for our research.
+4.2. Experiment 2. Sequential Recommendation
+4.2.1. Datasets and Setup
+In this study, our objective is to understand the generalizability of the findings from the previous study and address aspects that were not studied previously.
+In particular, Experiment 2 focuses on the more practical task of sequential recommendation and the comparison of zero-shot vs.
+few-shot in-context learning (ICL), which we evaluate on datasets of different domains (movies and music).
+The choice of sequential recommendation, rather than classical top-$k$, was motivated by the research carried out in (Hou et al., 2023).
+Therefore, the primary goal in Experiment 2 is not to assess the quality of different prompt scenarios (as carried out in Section 4.1), rather to evaluate the effectiveness of zero-shot vs.
+few-shot ICL on the accuracy vs.
+item fairness of RS.
+We focus our attention on two main datasets from the movie and music domains: MovieLens-1M and LastFM-1K.
+These datasets are widely recognized in the academic community, and provide a basis for evaluating our models in terms of fairness across different types of media consumption data.
+Here, we opt to use a sub-sample of users instead of all, to speed up experiments.
+Initially, we randomly select a subset of 80 users who exhibit a moderate level of interaction within the datasets.
+This allows us to handle the data efficiently while ensuring that the users selected have enough interactions to inform the training process.
+The data for these users is divided into training and test sets by sorting their interactions over time and splitting them so that 80% of a user interactions are used for training, with the remaining 20% held out for testing.
+This method respects the chronological order of interactions, thereby simulating a realistic scenario where a model can only learn from past data to make predictions about future user behavior.
+Table 6. Statistics of the final datasets used in the sequential ICL task.
+Dataset —U— —I— —R— $\frac{R}{U}$ $\frac{R}{I}$ Density Item Gini User Gini
+LastFM-1K 80 5,500 277,607 3,470.09 50.47 98.6% 0.697 0.621
+MovieLens 1M 80 3,900 20,987 262.34 5.38 99.2% 0.745 0.657
+4.2.2. Baselines.
+In this experiment, we do not compare the built RecLLMs with traditional sequential models, as this would require a separate and labor-intensive set of experiments, which is beyond the scope of this study.
+Instead, we compare the fairness and accuracy of different versions of RecLLMs against each other.
+Additionally, in one part of the experiment, we explore the impact of the presence or absence of user demographics in the prompts on recommendation fairness and quality.
+Since CF baselines typically do not use such demographic information, we exclude them from this comparison.
+4.2.3. Evaluation metrics
+The first experiment employs a diverse range of metrics to assess the performance of our recommendation system.
+These metrics are categorized into two primary classes: accuracy metrics, and item fairness, and catalogue coverage.
+Accuracy Metrics.
+For accuracy, we choose simpler metrics since the goal is not to compare them with strong CF baselines.
+We use the following two metrics to measure the quality of the top-$k$ recommendations in sequential setting:
+  • Hit Rate (HR@k): Measures the proportion of users for whom at least one of the top-$k$ recommended items is relevant. It is calculated as:
+[…]
+    where $R_{Q}$ represents the set of recommendations of the ranker in question and $\mathcal{G}$ represents the set of items in the ground truth.
+  • Average Rank (AverageRank@K): Measures the average position of relevant items in the recommendation list. It is calculated as:
+[…]
+    where $\text{rank}_{R_{Q}}(i)$ is the rank of item $i$ in the recommendation list $R_{Q}$.
+High scores in these metrics indicate more accurate and relevant recommendations.
+Item fairness, catalouge Coverage and long-tail distribution.
+Similar to the previous experiment, we use the Gini Index, HHI, and Entropy to measure item fairness.
+Additionally, we compute catalogue coverage, and display the long-tail distribution to visualize the concentration bias of recommendated items.
+4.3. Parsing Output Recommendations and Finding Closest Match Items
+Since language models (LLMs) naturally produce text-based outputs, we need a systematic approach to parse these outputs and map them to our existing dataset items.
+This ensures that when the RecLLM generates recommendations for which we do not find an exact match in the dataset, we can associate them with the most similar existing items, maintaining the integrity and relevance of our recommendation system.
+  • We start by utilizing the OpenAI GPT-3.5 API itself to parse the raw text of the recommendations.
+We designed a LLM_parsing function to extract the recommended songs and artists (or movies) from the text provided by the language model.
+For instance, we specify a system message to format the output in a structured way, ensuring that each recommendation follows the format: “Song Name by Artist.” This structured output helps the subsequent matching process.
+The parsed recommendations are then collected, which contain cleaned and structured text.
+  • Once the recommendations are parsed, the next step involves finding the closest matches to these recommendations within our dataset.
+This is achieved using the find_closest_match function, which employs the difflib.SequenceMatcher from the Python standard library.
+This tool calculates similarity scores between the recommended song and artist names and the existing items in our dataset.
+For each recommendation, the function compares it against all items in the dataset, computing a combined score based on the similarity of both the track name and the artist name.
+  • A threshold is set to ensure that only matches with a high enough similarity score are considered valid.
+If the highest score exceeds this threshold, the corresponding item ID from the dataset is selected as the closest match.
+This method ensures robustness by preventing low-similarity matches from being accepted, thus maintaining the relevance and accuracy of the recommendations.
+In summary, the implementation of these methods leverages essential Python tools and libraries, openai for API interaction, and difflib for sequence matching.
+These tools provide a robust and efficient framework for parsing and matching recommendations, ensuring that the RecLLM system can be effectively evaluated even when it generates items not originally present in the dataset.
+5. Results and Discussion
+In this section, we present the results and discussion for two experiments. We formulate hypotheses for each part, then showcase the results and provide analysis. Finally, we answer the hypotheses.
+5.1. Experiment 1. Zero-Shot Top-k Recommendation
+Through experiment 1, conducted in a zero-shot setting for the classical top-$k$ recommendation, we aim to answer the following experimental research questions.
+Details about Experiment 1 can be found in Section 3.1 and Section 4.1.
+We evaluate the average performance (Avg Perf.) in terms of NDCG and Recall across different conditions/models.
+We compare regular performance values (average) with bootstrap averages to gain deeper insights.
+The bootstrap method involves repeated resampling and offers a more robust understanding of performance variability.
+Table 7 illustrates this comparison.
+While the regular average provides a single estimate, the bootstrap mean extends this by offering confidence intervals, indicating the range within which the true performance measure is likely to fall.
+Experimental Research Questions RQ1.
+How does the incorporation of various goal-oriented prompts impact the accuracy (personalization) of GPT-based model in comparison to CF baselines?
+(cf.
+Section 5.1.1) RQ2.
+What is the stability and consistency of the personalization performance metrics (e.g., NDCG, Recall) for GPT-based recommendation systems across multiple runs?
+How does this variability compare to CF baselines?
+(cf.
+Section 5.1.2) RQ3.
+How do GPT-based recommendation systems compare to CF baselines in terms of provider fairness?
+Can the inclusion of a “Fair Recommender” system role might mitigate item unfairness and enhance diversity?
+(cf.
+Sections 5.1.3) RQ4.
+Do GPT-based recommendation systems exhibit a bias towards recommending newer or older movies, or certain genre tendancy, compared to CF baselines, and how does the temporal freshness of recommendations hold?
+How does the inclusion of explicit user ratings in prompts affect the personalization performance?
+(cf.
+Sections 5.1.4).
+RQ5.
+How do different algorithmic in-context learning (ICL) strategies, — in particular zero-shot vs.
+few-shot learning – impact the quality and biases of RecLLMs?
+What are the specific prompt design aspects that may influence this towards better or worse performance?
+(cf.
+Section 5.2) RQ6.
+What are the economic and practical implications of using GPT-based models for recommendation systems, specifically in terms of inference costs and latency issues?
+(cf.
+Section 5.3)
+5.1.1. Personalization of Recommendations
 This section focuses on the personalization performance measured by recommendation top-$K$ accuracy.
-Results can be found in Table 5, in which each generative scenario (i.e., using ChatGPT) could be found on the upper tabs while the result of CF baselines is on the bottom.
-It should be noted that, given the extensive nature of these experiments, not every possible scenario combination was tested independently; rather, a select subset of scenarios was examined.
-We will analyze the results by evaluating the average performance (abbreviated as Avg Perf.) in terms of NDCG and Recall across different conditions/models.
-These conditions include exploring the impact of (i) the system’s role: No Role (R0), Standard Recommender (R1), and Fairness-aware Recommender (R2), and (ii) the presence (E1) or absence (E0) of a fairness statement in the LLM prompts.
-Table 5. Table of the performance result measured in terms of recommendation Top-10 accuracy.
-“System” Role: No (R0), Emphasis: No (E0)
-Avg Perf. Bootstrap Mean Bootstrap Conf
+Results are presented in Table 7, with each generative scenario (using ChatGPT) in the upper sections and CF baselines in the bottom section.
+Not every possible scenario combination was tested; rather, a select subset was examined.
+Hypotheses
+To investigate the effectiveness of different recommendation strategies, we propose the following hypotheses:
+  • H1.
+Incorporating goal-oriented prompts, including personalization-based, beyond-accuracy, and reasoning-based prompts (as detailed in scenarios S1 to S7 in Section 3.1.1), with different system assignment roles, results in substantially varied performance outcomes in RecLLMs;
+  • H2. In terms of performance, GPT-based recommenders in zero-shot setting can provide comparable performance compared to collaborative filtering (CF) baselines.
+Discussion.
+Our evaluation highlights several findings regarding the personalization of recommendations using RecLLM.
+  • First, there is a clear indication that incorporating the “system role” improves the performance of RecLLM, and this consistency is observed across most scenarios.
+For example, for a randomly chosen scenario S1, in Tab1 and Tab2 (with no system role assignment), the NDCG performances are 0.008803 and 0.010431, respectively.
+In contrast, Tab3 and Tab4 (with system role assignments, either as “act as a recommender system” or “act as a fair recommender system”) show performance improvements to 0.013007 and 0.014268, respectively.
+This represents roughly a 45% improvement.
+In some scenarios, the improvement reaches up to 100%.
+It becomes thus evident that the system role plays a crucial part in optimizing the effectiveness of recommendation systems.
+  • Second, the performance varied considerably across different scenarios (S1 to S7), indicating that the context and nature of the prompts affect recommendation outcomes.
+For example, in Tab4, the “Simple” prompt scenario (S1) achieved an NDCG of 0.014, while the “Diversify” prompt (S4) had a lower NDCG of 0.002.
+The best-performing methods were simple, reasoning-based (COT and motivated reasoning), providing relatively better accuracy.
+Thus, based on these results, we may conclude that requesting ChatGPT to diversify recommendation or to provide novel recommendations significantly reduces accuracy, which should be avoided or carefully considered.
+Overall these variations underscore the importance of selecting appropriate prompts based on the desired recommendation attributes.
+  • Finally, comparing RecLLMs with baselines, our results reveal that RecLLM, even in zero-shot settings, generally performs lower than CF baselines, though in certain contexts, this difference shrinks considerably.
+For instance, the average NDCG of CF models such as BPR-MF and ItemKNN were 0.04776 and 0.04905, respectively.
+In comparison, the best-performing RecLLM prompt achieved an NDCG of 0.014, which is approximately 70% lower.
+Table 7. Performance results measured in terms of recommendation Top-10 accuracy.
+Tab 1 - System Role: No (R0), Emphasis: No (E0)
+Prompt: Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy.
+System Role: -
+Avg Perf. Bootstrap Mean Bootstrap Conf.
 Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
@@
 S6. Motivate Reasoning 0.010152 0.010906 0.010073 0.01098 (0.007446, 0.013124) (0.007511, 0.0148920)
-S7. Chain-of-thought (COT) 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
-“System” Role: No (R0), Emphasis: Fair (E1)
-Avg Perf. Avg. Bootstrap Bootstrap Cnf
+S7. COT 0.006573 0.005848 0.006554 0.005821 (0.004513, 0.008940) (0.003853, 0.008015)
+Tab 2 - System Role: No (R0), Emphasis: Fair (E1)
+Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend … .
+System Role: -
+Avg Perf. Avg. Bootstrap Bootstrap Conf.
 Type Model NDCG Recall NDCG Recall NDCG Recall
@@
 S6. Motivate Reasoning 0.004315 0.005097 0.004302 0.005031 (0.002246, 0.006819) (0.002668, 0.007909)
-S7. Chain-of-thought (COT) 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
-“System” Role: Normal Recommender (R1), Emphasis: Fair (E1)
-Avg Perf. Avg. Bootstrap Bootstrap Conf
+S7. COT 0.007207 0.009344 0.007177 0.009373 (0.004493, 0.010162) (0.005808, 0.013662)
+Tab 3 - System Role: Normal Recommender (R1), Emphasis: Fair (E1)
+Prompt: Based on these movies: {user_movies_string}, recommend 10 movies that the user will likely enjoy.
+System Role: Act as a fair recommender system balancing between popular and less-known movies to ensure provider fairness.
+Avg Perf. Avg. Bootstrap Bootstrap Conf.
 Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
@@
 S6. Motivate Reasoning 0.009379 0.00957 0.009447 0.009496 (0.006160, 0.012722) (0.006056, 0.013061)
-S7. Chain-of-thought (COT) 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
-“System” Role: Fair Recommender (R2), Emphasis: No (E0)
-Avg Perf. Avg. Bootstrap Bootstrap Cnf
+S7. COT 0.013571 0.012887 0.013622 0.012941 (0.009836, 0.018061) (0.009306, 0.017057)
+Tab 4 - System Role: Fair Recommender (R2), Emphasis: No (E0)
+Prompt: Ensure a fair representation of both popular and less-known movies. Based on these movies: {user_movies_string}, recommend … .
+System Role: Given a user, act as recommender system.
+Avg Perf. Avg. Bootstrap Bootstrap Conf.
 Type Model NDCG Recall NDCG Recall NDCG Conf. Recall Conf.
@@
 S6. Motivate Reasoning 0.012354 0.011968 0.012407 0.011929 (0.008901, 0.016671) (0.008133, 0.016052)
-S7. Chain-of-thought (COT) 0.013212 0.012862 0.013162 0.012908 (0.009972, 0.017185) (0.009256, 0.016893)
+S7. COT 0.013212 0.012862 0.013162 0.012908 (0.009972, 0.017185) (0.009256, 0.016893)
 CF Baselines
-Discriminative BPR-MF 0.047762 0.060699 0.047742 0.060537 (0.040465, 0.054958) (0.050181, 0.071453)
+Discriminative BPR-MF 0.04776 0.060699 0.047742 0.060537 (0.040465, 0.054958) (0.050181, 0.071453)
 ItemKNN 0.04905 0.065558 0.048995 0.065367 (0.041122, 0.056398) (0.054682, 0.076526)
@@
 TopPop 0.030556 0.032327 0.030539 0.032127 (0.024253, 0.036734) (0.025213, 0.039704)
-Additionally, a comparison between the regular performance values (obtained using average across values) and bootstrap averages of model performances provides deeper insights.
-The bootstrap method, which involves repeatedly resampling with replacement, offers a more robust understanding of the variability in performance measures.
-Table 5 illustrates this comparison.
-For instance, while the regular average provides a single estimate of NDCG and Recall, the bootstrap mean extends this by offering confidence intervals.
-These confidence intervals, represented as Bootstrap Conf.
-in the table, indicate the range within which the true performance measure is likely to fall with a certain level of confidence.
-We outline below the hypotheses regarding personalization that are central to this section:
-Hypothesis 1 (H1): The incorporation of a variety of goal-oriented prompts (as detailed in scenarios S1 to S7 in Section 3.1) across diverse settings (e.g., involving different system assignment roles) is anticipated to yield substantially varied performance outcomes in RecLLM.
-Hypothesis 2 (H2): GPT-based recommenders, particularly in a zero-shot setting, could achieve results comparable to CF baselines in specific scenarios.
-This is more probable in situations where the system is encouraged to engage in motivated reasoning and/or adopt COT strategies.
-This hypothesis is founded on observations that such strategies have demonstrated competitive performance in previous ML tasks.
-The Impact of Prompt Structure and Setting
-In the uppermost tab of Table 5, which presents the settings designated as No Role (R0) for the system role, the results are typically less favorable.
-Specifically, the prompt scenarios Simple (S1) and Motivated Reasoning (S5) demonstrate better performance within their respective segments.
-Scenario S1 under No Role (R0) reports NDCG and Recall values of 0.008803 and 0.00994, respectively.
-In comparison, S5 shows slightly improved metrics, with an NDCG of 0.010152 and a Recall of 0.010906.
-Overall, these results are the poorest when compared to other settings utilizing LLM-promoted prompts.
-Contrastingly, the data in tabs 3 and 4 indicate that when roles are delineated as either Normal Recommender (R1) or Fair Recommender (R2), there is a notable enhancement in both NDCG and Recall metrics.
-It is of particular note that under these roles (R1 and R2), scenarios such as simple (S1), chain-of-thought (COT) (S7), and Motivated Reasoning (S6) markedly exceed the performance metrics of their No Role (R0) counterparts.
-The most pronounced performance increase is evident in the fourth tab, where models reach their optimal enhancements.
-The simple model (S1) advances from 0.008803 to 0.014268 in NDCG (approximately a 62% increase) and from 0.00994 to 0.015796 in Recall (approximately a 59% increase).
-The Motivated Reasoning (S6) and COT (S7) models also exhibit substantial gains, with S6 escalating from 0.010152 to 0.012354 in NDCG (21.7% increase) and from 0.010906 to 0.011968 in Recall (9.7% increase), and S7 rising from 0.006573 to 0.013212 in NDCG (101% increase) and from 0.005848 to 0.012862 in Recall (120% increase).
-This trend is similarly observed in the Normal Recommender, Emphasis: Fair (R2/E1) configuration.
-These results emphasize the significant advantages of integrating system roles into recommendation systems.
-Bootstrap values in the R2-E0 setup further affirm the consistency and reliability of these improvements.
-Both the ‘simple’ and ‘COT’ models exhibit stable performance ranges across different roles and conditions.
-Comparison with CF Baselines
-When evaluating the effectiveness of GPT-based recommendation models compared to CF baselines, it is apparent that both the NDCG and the recall values for the GPT models are generally lower than those for the CF baselines.
-This suggests that, while GPT-based models have made significant strides, especially when designated with specific system roles such as “Normal Recommender” (R1) or “Fair Recommender” (R2), they still lag behind CF models in terms of both ranking relevant items at the top (NDCG) and retrieving a broader set of relevant items.
-For example, even with the “Fair Recommender” (R2) role applied, the “Simple” (S1) GPT-based model achieves an NDCG of 0.014268 and a recall of 0.015796, which are enhancements of the scenario “No Role” (R0) but still do not surpass the CF baseline models such as “BPR-MF,” “ItemKNN,” “NCF,” and “LightGCN,” that demonstrate higher efficacy in both metrics.
-The “TopPop” model, while having the lowest scores among the CF baselines, still presents an NDCG and Recall that are competitive with the improved GPT-based models.
-In summary, some noteworthy best and worst performances can be found below:
-  • For the GPT-based models under the “Fair Recommender” (R2) role with no fairness emphasis (E0):
-    • The Simple (S1) scenario achieves the best NDCG of 0.014268 and a Recall of 0.015796.
-    • The Diversify (S4) scenario shows the worst NDCG of 0.002844 and a Recall of 0.004901.
-  • For the CF baselines:
-    • The ItemKNN model has the highest performance with an NDCG of 0.04905 and a Recall of 0.06558.
-    • The TopPop model displays the lowest performance among the CF baselines with an NDCG of 0.030556 and a Recall of 0.032327.
-Despite the clear impact of prompts on the performance of RecLLM, it can be noted that Top-$k$ recommendation performance is not on par with traditional CF baselines.
-There are multiple strategies to address this issue, including enhancing prompt clarity and transitioning to few-shot learning.
-However, we posit that a significant factor contributing to this discrepancy is the possibility that many movies recommended by the GPT-recommender might not be present in the dataset.
-Therefore, by constraining GPT to select items only from the test item set, we believe that this could markedly improve the accuracy and relevance of the recommendations.
-H1: Supported.
-The introduction of specific roles (e.g., Normal and Fair Recommender) in recommendation systems significantly enhances the performance of generative models.
-This is evidenced by substantial increases in NDCG and Recall metrics when comparing the No Role scenario (R0) with Role-Based scenarios (R1, R2).
-It is worth noting in terms of which prompt scenario is effective, the Simple (S1) scenario, suggesting “Recommend 10 movies that the user will likely enjoy,” appears to provide the most effective approach for achieving the best top-$K$ recommendation accuracy.
-In other scenarios, only the reasoning scenario, such as COT (S7) matches this performance.
-(H2): Partially Supported.
-The evidence does not support Hypothesis 2 (H2), which posited that GPT-based recommenders could match CF baselines in certain settings.
-While GPT models, particularly with system roles and strategies scuh as Simple, ‘Motivated Reasoning’ and ‘COT’, show enhanced NDCG, Recall, they generally do not reach the performance of CF baselines.
-It is worthwhile that these results indicate a trend towards achieving outcomes closer to those of Top-Pop with the refinement of prompt design, better rationalizing, or transitioning to a few-shot learning setting.
-5.2. Stability and Confidence of Personalization Metrics
-As stated in Section 3.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models.
+Answer to Hypotheses
+To address the proposed hypotheses, we present the following findings:
+  • H1. $\checkmark$ Supported – The integration of specific system roles (e.g., “act as a recommender,” or “act as fair recommender”) leads to considerable performance improvements in RecLLMs.
+  • H2. $\times$ Rejected – Performance of RecLLMs in zero-shot settings consistently and significantly lags behind CF baselines.
+5.1.2. Stability and Confidence of Personalization Metrics
+As stated in Section 3.1.2, we considered the possibility that external factors such as trending data or algorithmic updates could introduce randomness into responses of GPT-based models.
 To study the impact of such randomness, each experiment was conducted five times specifically for GPT-based models.
-It is important to note that prior to our experiments, we lacked specific knowledge about the performance of Chat-GPT prompts.
-Therefore, our experiments focused exclusively on the R0/E0 scenario, which demonstrated poorer performance compared to the CF baseline.
-Despite this, the variability of results across different runs is noticeable.
-Consequently, in the analysis of this section, we should focus on observing the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
-Hypotheses on Stability of the Top-$K$ Accuracy Metric
-Hypothesis 1 (H1): Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent NDCG scores.
+[^4] Note that prior to these experiments, we lacked specific knowledge about the performance of Chat-GPT prompts.
+Our experiments thus focused on a randomly selected case (R0/E0 scenario).
+In the analysis of this section, our goal is to observe the variability of GPT-based models across different runs, rather than comparing them to the CF baselines.
+Hypotheses
+To rigorously evaluate the robustness and consistency of GPT-based models under varying conditions, we propose the following hypothesis:
+  • H3. Running multiple iterations of GPT-based recommendation queries for a specific user, results in consistent recommendation quality.
+Discussion
 The graph in Figures 3 illustrates the NDCG and Recall scores for GPT-based recommendation models across five different runs.
@@
 For instance,
-  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std333standard deviation of approximately 0.000305.
+  • Simple (S1): The NDCG scores are [0.0088, 0.0091, 0.0095, 0.0094, 0.0095], with a std555standard deviation of approximately 0.000305.
   • Motivate reasoning (S6): The NDCG scores are [0.0101, 0.0097, 0.0093, 0.0094, 0.0095], with a std of approximately 0.000316.
@@
 The dynamic nature of data and algorithms in generative models means that updates and changes over time could lead to different behaviors and outputs.
-Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the model’s performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
-H1: Supported.
-The data presented in the NDCG graph aligns with the hypothesis.
-While there is some expected variability across multiple runs, it remains within an acceptable/predictable range.
-Confidence intervals overlap significantly for different runs of each GPT-based model.
-While this consistency across runs confirms the system’s reliability in providing personalized recommendations, it is crucial to note that these findings do not necessarily guarantee stability over longer time frames, such as weeks or months.
-Therefore, it is advisable to approach these results with caution when considering long-term deployment.
-5.3. Diversity and Provider Fairness
-This section examines the diversity and fairness of different recommender models.
-Fairness is assessed by measuring the equitable distribution of recommendations across all items, avoiding bias towards popular or specific subsets (Gini Index, HHI).
-Diversity is gauged through the variety and uniqueness of the recommendations, ensuring the inclusion of a broad spectrum of items (Entropy), refer to Section 4.2.
-These evaluations are crucial to understanding each model’s effectiveness in matching user preferences and its capacity to foster a balanced and varied recommendation environment.
-H1: Given their training on extensive internet data, GPT-based models are hypothesized to exhibit a greater popularity bias than CF models, potentially leading to lower diversity in recommendations.
-This will be assessed using the Gini Coefficient and HHI for fairness, and Entropy for diversity.
-H2: The inclusion of a Fair Recommender as a ”system” role in GPT-based systems is expected to mitigate popularity bias and enhance the diversity of the system’s output.
-For H1, upon examining the results presented, in Table 6 a notable observation is that the ItemKNN model consistently exhibits the highest recommendation fairness (achieving the lowest Gini Coefficient and highest Entropy) among all models.
-This, to conduct a focused discussion, we focused on the remaining machine learning-based CF models and GPT-based models, we will exclude ItemKNN from subsequent comparisons.
-Without considering ItemKNN, the general trend suggests that the top-performing GPT models (Simple, Motivate Reasoning, and Chain-of-thought (COT)) are somewhat comparable to the majority of ML-based CF models, except when compared to NGCF, which shows a notably better performance in terms of fairness (Gini Coefficient: 0.950845 vs.
-0.0982346, Entropy: 6.420996 vs.
-5.042821) and diversity.
-Intriguingly, methods designed to promote diversification and novelty in GPT case (Surprise and Diversify Recommendation) perform contrarily, evidenced by their increased Gini Coefficients (Surprise: 0.997906, Diversify Recommendation: 0.992349) and reduced Entropy (Surprise: 3.227737, Diversify Recommendation: 4.232139), which is almost the worst in the respective scenarios, suggesting that these strategies may not effectively align with the objectives of reducing bias and enhancing diversity.
-Therefore, upon excluding the exceptional performance of ItemKNN, which seems to leverage its distinct mechanism effectively in a dataset characterized by popularity bias, the GPT-based models’ levels of fairness and diversity appear comparable to those of CF models.
-The data suggests that GPT-based models generally match their CF counterparts, with exceptions such as NGCF which shows superior fairness.
-Interestingly, NGCF, despite being the least performing CF model in some aspects, still outperforms GPT-based models on fairness.
-When we turn our attention to the top-performing CF models, namely BPR and LightGCN, it is evident that they achieve their enhanced performance by concentrating user preferences on a narrower selection of items, thus limiting diversity—a trend less likely to be observed in GPT models.
-This is attributable to the GPT models’ design, which is not inherently influenced by the popularity bias present in our datasets.
-For H2, the adaptation of GPT-based models with a ‘Fair Recommender’ system role results in a significant improvement in fairness and diversity metrics.
-Excluding ItemKNN for its good performance, the top GPT models exhibit better fairness in the Fair Recommender scenario, as evidenced by a noticeable decrease in Gini Coefficient (Simple: from 0.982463 to 0.978925, Chain-of-thought (COT): from 0.986889 to 0.979313) and HHI, coupled with an increase in Entropy (Simple: from 5.042821 to 5.387465, Chain-of-thought (COT): from 4.619500 to 5.365294).
-This trend suggests that GPT-based models, when guided by the system role within the ChatGPT framework, demonstrate a promising capacity for controllability.
-Such adaptability could mitigate potential biases inherent in the models—a flexibility that may not be as pronounced in traditional CF models.
-However, specific strategies such as Diversification (S4) and Surprise (S5) integrated into GPT-based models do not uniformly follow this positive trend.
-For example, in the Diversify scenario, we observe a slight increase in the model’s Gini coefficient (from 0.992349 to 0.992603) and a marginal rise in HHI (from 0.034724 to 0.030010), which deviates from the expected reduction.
-Similarly, the Surprise scenario shows a rise in Gini Coefficient (from 0.997906 to 0.998365) and HHI (from 0.059857 to 0.067952), alongside a reduction in Entropy (from 3.227737 to 3.023948).
-These anomalies highlight that certain approaches within the Fair Recommender role may inadvertently intensify the biases they aim to alleviate.
-The broader implication of this study is that the system role in the ChatGPT framework, alongside certain prompt scenarios, has the potential to align or even surpass the fairness and diversity benchmarks set by classical CF models.
-In notable instances, the bias is reduced, as indicated by lower Gini and HHI values, and increased Entropy.
-This indicates a potentially significant step forward in the controllability of GPT-based models to prevent their biases, an attribute that arguably gives them an edge over CF models in managing the diversity and fairness of recommendations.
-Answer to H1: Partially supported.
-The results do not uniformly support H1.
-While certain GPT-based models do exhibit higher popularity bias as indicated by the Gini Coefficient and HHI compared to some CF models, this is not a consistent finding across all models.
-For example, models such as NGCF show superior fairness indicators compared to some GPT-based models.
-However, when excluding the outlier performance of ItemKNN, GPT-based models are generally on par with CF models regarding fairness and diversity.
-Therefore, the hypothesis that GPT-based models inherently lead to lower diversity due to popularity bias is only partially supported.
-Answer to H2: Supported.
-The evidence supports H2 to a significant extent.
-When GPT-based models adopt a Fair Recommender role, there is a noticeable improvement in fairness and diversity metrics, such as a decrease in Gini Coefficient and HHI, along with an increase in Entropy.
-This indicates that GPT-based models have an inherent flexibility that, when leveraged with the right system role, can effectively mitigate popularity bias and enhance the diversity of recommendations.
-In both instances, it is critical to acknowledge that specific strategies, namely “diversification” and “surprise,” do not conform to their anticipated effectiveness—in some cases, they produce outcomes contrary to expectations.
-This observation implies that certain types of prompts may inadvertently introduce complexity, leading to a confounding of the system’s performance.
-Table 6. Diversity and provider fairness of the experimented models
+Nonetheless, caution is required when generalizing these results, and continuous monitoring and periodic re-evaluation of the performance over extended periods are recommended to ensure that the insights remain relevant and accurate.
+Answer to hypotheses
+To evaluate the validity of our proposed hypotheses, we present detailed findings below, supported by statistical data and analyses:
+  • H3.
+$\checkmark$ Supported – This hypothesis, which anticipated consistent performance across different GPT-based model runs, is supported.
+The data presented in the NDCG graph aligns with this expectation, showing only minor variability across multiple runs.
+Confidence intervals overlap significantly for different runs of each GPT-based model, which confirms the reliability of the system in providing personalized recommendations.
+However, it is important to consider these results with caution for long-term deployment as they do not necessarily guarantee stability over extended periods such as weeks or months.
+5.1.3. Item Fairness
+This section explores the item fairness metrics of different recommender models, crucial for evaluating their effectiveness in providing balanced and varied recommendations.
+Hypotheses
+To ascertain the impact of different model strategies on recommendation fairness, we carefully analyzed the outcomes in the context of our formulated hypotheses:
+  • H4: Due to their training on large-scale internet data, GPT-based models are hypothesized to exhibit lower item fairness than classical CF models.
+However, asking the system to produce more diverse or additional recommendations (scenarios S3 and S4) is expected to increase/improve item fairness.
+  • H5: Integrating a “Fair Recommender” as a system role in GPT-based systems is expected to enhance item fairness.
+Table 8. Provider fairness of the tested models in Experiment 1
 Model Normal Recommender Fair Recommender
-Gini Coefficient HHI Entropy Gini Coefficient HHI Entropy
+[…]
 Simple 0.982463 0.017204 5.042821 0.978925 0.010899 5.387465
@@
 TopPop 0.994859 0.020000 3.912023 0.994859 0.020000 3.912023
-5.4. Temporal Freshness Bias in Movie Recommendations
-In this section, we study the recency and freshness of the movie recommendations provided by various GPT-based and CF models.
-The investigation focuses on whether the models demonstrate a bias toward recommending newer or older films.
-By examining the mean and median release years of the movies suggested by each model, we can gauge their tendency toward recency.
+Cyan shows the best performing methods.
+Green shows good performing methods (relative to others).
+Yellow shows lower performance models.
+Results and Discussion
+We now address the above research hypotheses, using Table 8 to inform our discussion.
+We primarily use the Gini coefficient and entropy for this analysis, noting that HHI shows somewhat similar performance to the Gini coefficient.
+Overall, by examining the Gini indices and entropy values, we observe that two CF models, Item-KNN and NGCF, provide the best performances in terms of item fairness.
+Among the RecLLM models, the Genre-focused, Motivated Reasoning, and Chain-of-thought scenarios exhibit notable item fairness after the CF baselines.
+Summarizing the best-performing models, we have: $ItemKNN/NGCF>Genre/Motivate/COT>rest$.
+One notable and interesting result from the table is that asking GPT to produce novel or diverse recommendations has little or no significant impact on improving its recommendability.
+The entropy generally increases for RecLLM models when switching to Fair recommenders, indicating higher novelty in recommendations.
+This pattern is not observed in the CF models, which show no change in their metrics.
+To address H4, we find that GPT-based models generally exhibit lower item fairness compared to classical CF models, as hypothesized.
+However, contrary to the expectation, asking the system to produce more diverse or additional recommendations did not significantly increase item fairness.
+For H5, integrating a “Fair Recommender” role in GPT-based systems did enhance recommendation fairness, as indicated by the increased entropy in the RecLLM models.
+This suggests a mitigation of item unfairness and an enhancement of recommendation diversity.
+Answer to Hypotheses
+We can answer to the above hypotheses as following:
+  • H4: $\circ$ Partially Supported — While classical CF models such as Item-KNN and NGCF exhibit higher fairness compared to RecLLM models, the GPT-based models named ”Genre-focused” and ”Motivated Reasoning” showed improvements in fairness.
+This hypothesis is partially supported as diversification strategies and requests for more novel recommendations did not increase item fairness in RecLLM models.
+  • H5: $\checkmark$ Supported — Implementing a ”Fair Recommender” role within GPT-based systems leads to noticeable improvements in fairness.
+This supports the hypothesis that transitioning from normal to fair recommender systems enhances fairness, demonstrating the controllability in RecLLM models, particularly in terms of diversity and novelty.
+5.1.4. Analysis of Temporal Freshness and Genre Bias in Recommender Systems
+In this unified analysis, we explore the tendencies of various GPT-based and CF models in terms of recency and genre preferences in movie recommendations.
+Our study aims to discern whether these models demonstrate a bias towards recommending newer or older films and if they exhibit specific genre preferences that could illuminate performance differences between model types.
+Research Hypotheses
+We provide the following statement of research hypotheses:
+  • H6.
+We hypothesize that GPT-based models, given their extensive training on diverse and recent datasets, are more likely to recommend newer films compared to classical CF models, which may favor older, well-established films.
+  • H7.
+We expect that GPT-based models will show a broader genre distribution in their recommendations due to their training on diverse data sources, compared to CF models, which may adhere more strictly to users’ historical genre preferences.
 Figure 4. Distribution of Movie Release Years as recommended by different models
-Table 7 presents the statistics that inform our understanding of each model’s tendency towards recommending newer or older films.
-Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007.
-This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense.
-Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself.
-In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films.
-This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired.
-The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
 Table 9. Model Statistics Reflecting Recency Bias
@@
 Pop 1995 5.65
-Answer to H1: Contrary to the hypothesis, the results reveals that GPT-based models, especially ‘Diversify’ (Median Year: 2007) and ‘Surprise’ (Median Year: 2006), tend to recommend significantly newer movies than CF models.
-CF models such as BPR-MF (Median Year: 1995) and Pop (Median Year: 1995) generally suggest older films, as reflected in the lower mean and median years.
-The larger standard deviation in years for GPT-based models like ‘Surprise’ (Std Year: 10.62) also indicates a broader temporal range in their recommendations.
-This suggests that GPT-based models, perhaps due to their ability to capture more recent trends and discussions from diverse datasets, are capable of recommending more recent content than CF models, thus refuting the initial hypothesis.
-5.5. Analysis of Genre Bias/Affinity in Recommender Systems
-The goal of this section is to examine the genre biases present in movie recommendations provided by Collaborative Filtering (CF) models and GPT-based models.
-By comparing the prevalence of genre terms in the recommendations, we aim to understand if a particular model, or class of models, has a specific genre affinity (or bias) which could illuminate the performance differences between these models.
-H1: We hypothesize that the differences in recommendation performance between CF and GPT-based models may be attributed to their respective genre preferences.
-CF models are expected to closely adhere to genre affinity in users’ profiles, while GPT models may exhibit a predominant recommendation of certain genres over others.
+Results and Discussions
+Table 9 presents the statistics that inform our understanding of each model tendency towards recommending newer or older films.
+Notably, the GPT-based models ‘Surprise’ and ‘Diversify’ push the boundaries towards recommending more recent items, with median release years of 2006 and 2007.
+This indicates a clear orientation towards newer movies, possibly interpreting ‘surprise’ and ‘diversification’ in their temporal sense.
+Since GPT models do not possess inherent information about the novelty of content in terms of user interactions, they likely interpret these directives based on the freshness of the content itself.
+In contrast, the CF models tend to recommend older films, median release years consistently at 1995, reflecting a stronger inclination towards well-established, historically popular films.
+This distinction suggests that GPT-based models might be leveraging their extensive training datasets to prioritize more recent films, which can be beneficial in scenarios where up-to-date recommendations are desired.
+The broader temporal range of recommendations from GPT-based models, as evidenced by the standard deviation in the year of movies recommended (Surprise: Std Year 10.62, Diversify: Std Year 11.76), further underscores their flexibility in catering to diverse temporal tastes.
 Upon analyzing the word clouds, it is evident that CF models such as BPR-MF, LightGCN, and VAE predominantly recommend genres such as “Action,” “Adventure,” “Thriller,” and “Sci-Fi.” These genres appear frequently and in larger font sizes, indicating their dominance in the recommendations made by these models.
@@
 Top model correspond to CF models (BPR-MF, LightGCN, RecVAE), while lower models include GPT-based recommenders (Simple, Diversity, COT).)
-Answer to H1: The analysis supports the hypothesis to a certain extent.
-While CF models tend to reinforce existing user preferences by frequently recommending popular genres such as “Action” and “Sci-Fi,” GPT-based models demonstrate a broader genre coverage in their recommendations.
-The presence of a more diverse set of genres like “Comedy,” “Drama,” and “Romance” in GPT-based recommendations suggests these models may be utilizing their extensive training on diverse data sources to introduce a wider variety of content.
-This could potentially lead to a discovery of new interests for users, though it may sometimes diverge from their established preferences.
-Such a tendency for broader genre representation by GPT-based models could be a contributing factor to their distinct performance characteristics compared to CF models.
-5.6. Implicit vs. explicit scenario
+  • H6.
+$\checkmark$ Supported – GPT-based models, particularly Diversify and Surprise, tend to recommend significantly newer movies than CF models.
+This finding supports the hypothesis, confirming that GPT-based models are effective in capturing recent trends and offering more temporally relevant recommendations.
+  • H7.
+$\circ$ Partially Supported – While CF models are found to predominantly recommend genres such as Action and Sci-Fi, GPT-based models demonstrate a broader genre distribution, including Comedy, Drama, and Romance.
+This diversity in GPT-based recommendations supports the hypothesis to a certain extent, indicating that these models are utilizing their extensive training on diverse datasets to introduce a wider variety of content.
+Implicit vs.
+explicit scenario.
 Analyzing the impact of including ratings in the prompt shows noteworthy results.
@@
 The confidence intervals for NDCG (0.009737 to 0.017012) and Recall (0.010261 to 0.019795) further emphasize its reliability.
-These results suggest that including ratings significantly enhances the recommendation system’s performance.
+These results suggest that including ratings significantly enhances the recommendation performance of the system.
 It provides a more personalized and accurate reflection of user preferences, which is a key aspect in recommendation systems.
@@
 This insight could be valuable in refining other scenarios and models, potentially leading to better overall system performance.
+5.2. Sequential Recommendations
+This section focuses on the performance of RecLLMs in sequential recommendation task, measured by recommendation top-$K$ accuracy, item fairness, and coverage.
+Results are presented in Table 10, with each generative scenario (using ChatGPT) in the upper sections and CF baselines in the bottom section.
+Hypotheses
+To investigate the effectiveness of different recommendation strategies, we propose the following hypotheses:
+  • H1. Incorporating sequential in-context learning (ICL) prompts, including ICL-1 and ICL-2, results in improved recommendation accuracy compared to zero-shot learning.
+  • H2. Sequential in-context learning (ICL) prompts will result in better item fairness and coverage compared to zero-shot learning.
+  • H3. A combination of improved accuracy, item fairness, and coverage will be observed in certain settings, highlighting the optimal recommendation strategy.
+Discussion.
+Our evaluation highlights several findings regarding the personalization of recommendations using RecLLM.
+  • Regarding H1, the results indicate that recommendation accuracy, measured by Hit Rate (HR) and NDCG, varies with different learning strategies.
+For the MovieLens dataset, without demographic information, the HR for zero-shot learning is 0.179, which decreases slightly for ICL-1 (0.154) and ICL-2 (0.150).
+In the LastFM dataset, zero-shot learning HR is 0.204, which decreases to 0.154 for both ICL-1 and ICL-2.
+    However, including demographic information, especially in the Age-Group category, improves accuracy in the movie domain.
+For instance, in the MovieLens dataset with age-group context, the HR for ICL-2 improves to 0.204, which is higher than zero-shot learning with such information (0.171) or zero-shot learning with no demographic information (0.179).
+Another clear trend is the impact of the Recency-Focused (Rec.-Freq) sampling in the context part on recommendation accuracy.
+When sensitive attributes are included, Rec-Freq and ICL-2 generally show better performance compared to the same age-category included in zero-shot learning.
+For example:
+    • In the MovieLens dataset with age-group context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.171 for zero-shot learning.
+    • In the MovieLens dataset with gender context, the HR for Rec-Freq ICL-2 is 0.175, compared to 0.158 for zero-shot learning.
+    • In the LastFM dataset with age-group context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.263 for zero-shot learning.
+    • In the LastFM dataset with gender context, the HR for Rec-Freq ICL-2 is 0.204, compared to 0.221 for zero-shot learning.
+    In terms of NDCG, however, results generally favor zero-shot learning scenarios across different combinations of demographic information and sampling strategies in both datasets.
+While there is a slight improvement in NDCG for the Rec-Freq ICL-2 scenario in the LastFM dataset compared to other scenarios (zero-shot or ICL-1), the differences are marginal.
+For example:
+    • In the LastFM dataset with age-group context, the NDCG for Rec-Freq ICL-2 is 0.60000, compared to 0.56250 for ICL-1, and 0.82500 for zero-shot learning.
+    • In the LastFM dataset with gender context, the NDCG for Rec-Freq ICL-2 is 0.65625, compared to 0.59375 for ICL-1, and 0.64375 for zero-shot learning.
+    These examples illustrate that incorporating sensitive attributes and employing the Rec-Freq sampling method in ICL-2 can enhance recommendation accuracy, although the improvements in NDCG are not as pronounced.
+  • Regarding H2, item fairness and coverage show mixed results across different settings.
+For item fairness, as measured by Gini and Entropy, zero-shot learning and ICL strategies show slight differences.
+For instance, in the MovieLens dataset with no demographic info, Gini for zero-shot learning is 0.57739 (lower is better), which increases to 0.66303 for ICL-1 and 0.65835 for ICL-2, indicating worse fairness.
+Entropy, where higher values are better, decreases from 3.47952 (zero-shot) to 2.87493 (ICL-1) and 2.88530 (ICL-2), suggesting that ICL prompts might lead to slightly worse item fairness.
+Coverage remains consistently low across all scenarios in the movie domain, and zero-shot learning provides considerably better coverage performance.
+For example:
+    • In the MovieLens dataset without demographic information, the coverage for zero-shot learning is 0.02621, compared to 0.01685 for ICL-1 and ICL-2.
+    • In the MovieLens dataset with gender context, the coverage for zero-shot learning is 0.02845, compared to 0.01498 for ICL-2.
+    In the music domain, results generally favor ICL-2. For example:
+    • In the LastFM dataset with age-group context, the coverage for Rec-Freq ICL-2 is 0.00210, compared to 0.00198 for zero-shot learning.
+    • In the LastFM dataset with gender context, the coverage for Rec-Freq ICL-2 is 0.00218, compared to 0.00211 for zero-shot learning.
+    Notably, including demographic information generally enhances fairness metrics.
+For example, with Gender context in MovieLens, the Gini index improves, and Entropy increases, indicating fairer item distribution.
+  • Regarding H3, a combination of improved accuracy, item fairness, and coverage is generally observed with the Recency-Focused approach in both zero-shot and ICL-2 scenarios.
+By visually examining the placements of highlighted values, we note that Recency-Focused zero-shot and ICL-2 tend to provide the best combined performance of accuracy, item fairness, and coverage.
+Additionally, revealing demographic information, particularly age-group (young/old) and gender, has a noticeable positive impact on the fairness and coverage of the methods, thereby enhancing the utility of the system from both consumer and producer perspective.
+Table 10. Recommendation alignment between ($\mathcal{R}_{m}$,  $\mathcal{R}_{m}^{a}$) based on Item Similarity $\beta_{item}$
+Attribute Type ICL MovieLens LastFM
+Acc. Item Fairness Cov. Acc. Item Fairness Cov.
+HR NDCG Gini Entropy - HR NDCG Gini Entropy -
+No Info Freq. 0-shot 0.179 0.53750 0.57739 3.47952 0.02621 0.204 0.66875 0.24882 4.89089 0.00195
+ICL-1 0.154 0.55000 0.66303 2.87493 0.01685 0.154 0.47500 0.24464 4.91536 0.00201
+ICL-2 0.150 0.40000 0.65835 2.88530 0.01685 0.154 0.44375 0.25137 4.92419 0.00201
+Rec.-Freq. 0-shot 0.208 0.65000 0.55075 3.66855 0.03033 0.225 0.65000 0.18229 5.07136 0.00220
+ICL-1 0.162 0.54375 0.60812 3.08940 0.01760 0.183 0.53750 0.20217 4.99446 0.00206
+ICL-2 0.175 0.56875 0.64473 2.92146 0.01610 0.208 0.58125 0.20439 5.00125 0.00205
+Gender Freq. 0-shot 0.146 0.58750 0.56064 3.60813 0.02845 0.196 0.68750 0.21299 5.00305 0.00211
+ICL-1 0.167 0.50625 0.67541 2.71790 0.01460 0.158 0.51875 0.26066 4.82855 0.00188
+ICL-2 0.171 0.43125 0.66145 2.76958 0.01498 0.154 0.42500 0.22654 4.99862 0.00215
+Rec.-Freq. 0-shot 0.158 0.55000 0.54779 3.68560 0.02958 0.221 0.64375 0.18254 5.08343 0.00223
+ICL-1 0.154 0.46250 0.64715 2.99814 0.01909 0.208 0.59375 0.19480 5.01322 0.00209
+ICL-2 0.175 0.50625 0.64839 2.73183 0.01310 0.204 0.65625 0.18622 5.06197 0.00218
+Age-Group Freq. 0-shot 0.154 0.70000 0.54193 3.72305 0.03145 0.183 0.56875 0.24498 4.91475 0.00198
+ICL-1 0.175 0.57500 0.67186 2.73694 0.01498 0.146 0.48750 0.24892 4.90102 0.00199
+ICL-2 0.150 0.45625 0.64637 2.87851 0.01572 0.146 0.41250 0.23557 4.97022 0.00211
+Rec.-Freq. 0-shot 0.171 0.66875 0.56433 3.63473 0.02995 0.263 0.82500 0.18816 5.07481 0.00223
+ICL-1 0.158 0.46875 0.63580 2.92408 0.01610 0.188 0.56250 0.20575 4.96456 0.00200
+ICL-2 0.204 0.60000 0.63038 2.93595 0.01498 0.204 0.60000 0.19985 5.01733 0.00210
+Cyan shows the best performing methods.
+Green shows good performing methods (relative to others).
+Yellow shows lower performance models.
+Answer to Hypotheses
+To address the proposed hypotheses, we present the following findings:
+  • H1.
+$\circ$ Partially Supported – The incorporation of sequential in-context learning prompts (ICL-1 and ICL-2) improves recommendation accuracy in certain contexts, particularly with demographic information (e.g., Age-Group in MovieLens), though not consistently across all settings.
+  • H2.
+$\times$ Rejected – Sequential in-context learning prompts do not consistently lead to better item fairness and coverage compared to zero-shot learning.
+In some cases, they worsen item fairness, as indicated by higher Gini indices and lower Entropy values.
+  • H3.
+$\checkmark$ Supported – Certain settings, particularly Recency-Focused ICL-2 with demographic information, show a balance between accuracy and item fairness, indicating potential optimal recommendation strategies.
+5.3. Economic and Practical Implications
+Integrating large language models (LLMs) like GPT-3.5-turbo into recommender systems involves several economic and practical considerations.
+Below, we discuss the inference costs, latency issues, and potential mitigation strategies.
+Inference Costs
+Using LLMs involves significant API costs. For example, the cost per API call can be calculated as:
+Cost per call=(tokens per $\text{Cost per call}=\left(\frac{\text{tokens per call}}{1000}\right)\times\alpha$
+where $\alpha$ is the cost per 1000 tokens. For a system with $n$ users and $m$ prompts, the total cost $C$ is:
+$C=n\times m\times\text{Cost per call}$
+For example, if an average recommendation call uses 1500 tokens, the cost per call is $0.03$.
+For 610 users with 7 prompts each, the total cost would be $128.10$.
+To mitigate these costs, we can explore strategies such as token optimization, batch processing, and selective use of LLMs.
+Our code implements these strategies by limiting the maximum tokens per response and grouping prompts to minimize the number of API calls.
+Latency
+Latency, the time taken to generate recommendations, affects user experience.
+Observed latencies ranged from a few to over ten seconds per call.
+Mitigation strategies include asynchronous processing to handle multiple requests concurrently, caching mechanisms to store common recommendations, and using faster, distilled versions of LLMs.
+Our code addresses latency by measuring and optimizing the time taken for each API call and incorporating mechanisms to batch and streamline requests.
+Additionally, practical considerations such as using cloud services for scalability, combining LLMs with traditional collaborative filtering models, and regularly evaluating performance metrics ensure the system remains efficient and cost-effective.
+Summary.
+Answers to Research Questions for Experiment 1 and 2 Answer to RQ1: The incorporation of various goal-oriented prompts impacts the accuracy of GPT-based models, with some prompts enhancing performance while others, like diversity-focused prompts, reducing it.
+However, GPT-based models generally lag behind CF baselines in zero-shot settings.
+Answer to RQ2: GPT-based recommendation systems show consistent personalization performance across multiple runs, with minor variability.
+This stability is comparable to CF baselines, indicating reliable behavior despite inherent randomness in individual runs.
+Answer to RQ3: GPT-based recommendation systems exhibit lower provider fairness compared to CF baselines.
+However, incorporating a “Fair Recommender” system role in GPT-based models improves fairness and diversity, demonstrating controllability and potential for mitigation of item unfairness.
+Answer to RQ4: GPT-based models tend to recommend newer movies and show a broader genre distribution compared to CF baselines, which prefer older, well-established films.
+Including explicit user ratings in prompts enhances personalization performance significantly.
+Answer to RQ5: Different ICL strategies impact RecLLM quality and biases variably, with zero-shot learning generally providing better accuracy, while ICL strategies, especially with demographic information, offer improvements in certain contexts.
+Prompt design plays a crucial role in these outcomes.
+Answer to RQ6: The economic and practical implications of using GPT-based models include significant inference costs and latency issues.
+Mitigation strategies like token optimization, batch processing, and using distilled models can help manage these challenges, ensuring efficient and cost-effective system performance.
 6. Conclusion and Future Directions
-This research presents a comprehensive analysis of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems.
-The study emphasizes the nuanced capabilities and inherent biases of these systems, highlighting their contrasting behaviors compared to traditional CF in the domain of movie recommendations.
-Our experimental findings underscore the significant impact of ‘prompt design’ strategies on recommendation quality, touching upon aspects such as accuracy, provider fairness, diversity, temporal stability, genre dominance, and freshness (recency).
-We discovered that role-based prompts (those using clear system roles such as you a Recommender, or you are a Fair recommender) enhance fairness and diversity, effectively mitigating popularity bias.
-Interestingly, GPT-based models, while not always paralleling the performance of CF baselines, exhibit a preference for recommending newer and more diverse movie genres.
-Moreover, our experiments reveal that including ratings in prompts appears to improve the personalization aspect of recommendations.
-GPT-based models also have a distinct inclination towards recommending newer, particularly those released post-2000.
-We hypothesize that a key reason for certain discrepancies in recommendations quality is that the GPT model might suggest movies not included in the dataset.
-By limiting the GPT model to choose only from a predefined set of test items, we anticipate a notable improvement in both the accuracy and relevance of its recommendations.
-Additionally, transitioning to a zero-shot scenario could significantly enhance personalization results in this context.
-In genral, we think that it important to balance personalization power of CF models with the broader exploratory nature of GPT-based RecLLMs, to ensure a diverse and engaging user experience.
-Key areas for future research include:
-  1. (1)
-    Refine Prompt Design Strategies and Dataset Constraints.
-This involves developing prompts that effectively balance user preferences with broader content exploration, while also constraining the system to select items only from within the dataset.
-This limitation is expected to improve the accuracy and relevance of recommendations.
-  2. (2)
-    Explore Few-shot Learning to Improve accuracy and Contextual Relevance.
-Few-shot learning involves exposing the model to a small number of examples from the target domain, which can enhance its ability to make more informed recommendations.
-  3. (3)
-    Integrate GPT with CF in a RAG framework.
-Aimed to create a hybrid approach that leverages the strengths of both methodologies.
-This combination could potentially lead to more robust and effective recommender systems.
-  4. (4)
-    Expand research to cross-domain applications and conduct long-term user engagement studies.
-This includes evaluating RecLLMs in various content domains (e.g., books, music, e-commerce) and studying user interaction over extended periods to gauge satisfaction, trust, and the evolving nature of user preferences.
-By focusing on these key areas, future research can significantly advance the field of recommender systems, particularly in leveraging the capabilities of large language models like GPT in diverse and dynamic recommendation scenarios.
+This research provides a comprehensive analysis of biases of Recommender Systems using Large Language Models (RecLLMs), specifically focusing on ChatGPT-based systems.
+Our study highlights the distinct capabilities and biases of these systems compared to traditional Collaborative Filtering (CF) methods within the domain of movie and music recommendations.
+The experimental results emphasize the significant impact of prompt design strategies on various aspects of recommendation quality, including accuracy, provider fairness, catalog coverage, diversity, temporal stability, genre dominance, and recency.
+In particular, we conducted two experiments using different datasets to evaluate the effectiveness of these strategies: classical top-$K$ recommendations and sequential in-context learning (ICL).
+Key contributions and findings of our work include:
+  • Enhanced Fairness and Diversity through Role-Based Prompts: Utilizing role-based prompts (e.g., “act as a recommender” or “act as a fair recommender”) significantly enhances fairness and diversity in recommendations, effectively mitigating item unfairness.
+  • Sequential In-Context Learning (ICL): Incorporating sequential in-context learning prompts (ICL-1 and ICL-2) shows improvements in recommendation accuracy in certain contexts, particularly with demographic information.
+For instance, in the MovieLens dataset with age-group context, the Hit Rate (HR) for ICL-2 improves compared to zero-shot learning.
+However, ICL prompts do not consistently lead to better item fairness and coverage compared to zero-shot learning.
+  • Improved Personalization with Ratings: Incorporating user ratings into prompts improves the personalization of recommendations, highlighting the potential for fine-tuning prompts to better reflect user preferences.
+  • Temporal Bias towards Recent Releases: GPT-based models exhibit a strong tendency to recommend newer movies, particularly those released post-2000, contrasting with the older preferences of CF models.
+This suggests that GPT models can capture recent trends better, making them more temporally relevant.
+  • Item Fairness and Diversity: While GPT-based models generally exhibit lower item fairness compared to classical CF models, scenarios involving fair recommender roles show improved fairness and diversity, as indicated by higher entropy values.
+  • Stability in Performance Metrics: Despite inherent randomness, GPT-based models show consistent performance across multiple runs, as evidenced by low standard deviations in metrics such as NDCG.
+This stability is crucial for reliable recommendation systems.
+Our research identifies several key areas for further exploration that can push forward the capabilities and effectiveness of recommendation systems leveraging large language models (RecLLMs).
+Expanding research to include various content domains such as books, music, and e-commerce, along with conducting long-term user engagement studies, will provide deeper insights into this research.
+Firstly, refining prompt design and enhancing generalization across various datasets and tasks is a crucial step to improving recommendation accuracy and relevance.
+Additionally, implementing more advanced few-shot learning techniques, such as carefully selecting which examples to use and how to present them, can enhance the contextual relevance and accuracy of recommendations.
+These examples could be chosen from similar users (peer users) to make the system behave more like CFg models, thereby combining the strengths of both approaches.
+Moreover, integrating GPT with collaborative filtering (CF) in a Retrieval-Augmented Generation (RAG) framework could leverage the strengths of both methodologies, leading to more robust and effective recommender systems.
+Future work should also address additional fairness dimensions, such as ensuring consumer-side fairness to prevent recommendations from disproportionately favoring certain user groups.
+Calibration, or aligning recommendations with user expectations to offer a balanced mix of familiar and novel items, is another critical aspect.
+Furthermore, ensuring counterfactual fairness, where recommendations remain fair even when user attributes or behaviors are altered, is also an essential and interesting future direction.
+We explicitly acknowledge the limitations of our current work in using a single LLM based on ChatGPT.
+Future studies should explore other large language models (LLMs) and compare their performance to further understand their potential and limitations in diverse recommendation scenarios.